### v3

# Directed schemes of ideals and cardinal characteristics, I: the meager additive ideal

Abstract.
We introduce the notion of directed scheme of ideals to characterize peculiar ideals on the reals, which comes from a formalization of the framework of Yorioka ideals for strong measure zero sets. We prove general theorems for directed schemes and propose a directed scheme $\vec{\mathcal{M}}=\{\mathcal{M}_{I}\colon I\in\mathbb{I}\}$ for the ideal $\mathcal{MA}$ of meager-additive sets of reals. This directed scheme does not only helps us to understand more the combinatorics of $\mathcal{MA}$ and its cardinal characteristics, but provides us new characterizations of the additivity and cofinality numbers of the meager ideal of the reals.
In addition, we display connections between the characteristics associated with $\mathcal{M}_{I}$ and other classical characteristics. Furthermore, we demonstrate the consistency of $\mathrm{cov}(\mathcal{N}\!\mathcal{A})<\mathfrak{c}$ and $\mathrm{cof}(\mathcal{MA})<\mathrm{non}(\mathcal{SN})$. The first one answers a question raised by the authors in [CMR25].
Key words and phrases:
Directed scheme of ideals, meager-additive ideal, cardinal characteristics, null-additive ideal, forcing models
2020 Mathematics Subject Classification:
03E17, 03E05, 03E35
The first author was partially supported by the Israel Science Foundation by grant 2320/23 (2023-2027); The second author was supported by the Grant-in-Aid for Scientific Research (C) 23K03198, Japan Society for the Promotion of Science; and the first and third authors were supported by the grant No. AP0010, Dirección de Tecnología e Investigación and Oficina de Internacionalización, Institución Universitaria Pascual Bravo
1. Introduction
Recent study of $\mathcal{SN}$, the ideal of strong measure zero subsets of $^{\omega}2$, and its cardinal characteristics, involves its characterization via the Yorioka ideals. These have been used to control de cofinality of $\mathcal{SN}$ in forcing models [Yor02, CM19, Car22, CMR22, CM25], as well as its additivity number [BCM25]. These ideals have helped to overcome challenges to control the combinatorics of $\mathcal{SN}$ and its cardinal characteristics, particularly in forcing models.
We aim to generalize the framework of Yorioka ideals to control de combinatorics of other very peculiar ideals. In this paper, we give special attention to the ideal of meager-additive subsets of $^{\omega}2$.
Notation 1.1.
Before proceeding, we first fix some notation.
  1. (1)
    Given a formula $\phi$, $\forall^{\infty}n<\omega\colon\phi$ means that all but finitely many natural numbers satisfy $\phi$; $\exists^{\infty}n<\omega\colon\phi$ means that infinitely many natural numbers satisfy $\phi$.
  2. (2)
    $^{\uparrow\omega}\omega$ denotes the set of all increasing functions in $^{\omega}\omega$.
  3. (3)
    For $\in{}^{\omega}\omega, xyx\leq^{*}y$ means that $\forall^{\infty}n<\omega\colon x(i)\leq y(i)$.
  4. (4)
    The identity function on $\omega$ is denoted by $\operatorname{id}$.
  5. (5)
    $\mathfrak{c}:=2^{\aleph_{0}}$; $\mathcal{N}$ and $\mathcal{M}$ denote the ideals of Lebesgue measure zero sets and of meager sets in the Cantor space $^{\omega}2$, respectively. Let $\mathcal{E}$ be the $\sigma$-ideal generated by the closed measure zero subsets of $^{\omega}2$. It is well-known that $\mathcal{E}\subseteq\mathcal{N}\cap\mathcal{M}$. Even more, it was proved that $\mathcal{E}$ is a proper subideal of $\mathcal{N}\cap\mathcal{M}$ (see [BJ95, Lemma 2.6.1]).
  6. (6)
    A preorder is a pair $\langle P,\leq\rangle$ where $P$ is a set and $\leq$ is a reflexive and transitive relation on $P$.
  7. (7)
    A preorder $\langle S,\leq\rangle$ is directed if, for $s,s^{\prime}\in S$, there is some $t\in S$ such that $s\leq t$ and $s^{\prime}\leq t$.
  8. (8)
    $\mathbb{I}$ denotes the set of partitions $I=\langle I_{n}\colon n<\omega\rangle$ of $\omega$ into consecutive non-empty intervals. We consider the directer preorder on $\mathbb{I}$ defined by $I\sqsubseteq J$ iff $\forall^{\infty}n<\omega\,\exists k<\omega\colon I_{k}\subseteq J_{n}$.
Given an ideal $\mathcal{I}$ of subsets of $X$ such that $\{x\}\in\mathcal{I}$ for all $x\in X$, the cardinal characteristics associated with $\mathcal{I}$ are defined by
$\operatorname{add}(\mathcal{I})$ $:=\min\{|\mathcal{J}|\colon\mathcal{J}\subseteq\mathcal{I}\text{\ and }\bigcup\mathcal{J}\notin\mathcal{I}\};$
$\operatorname{cov}(\mathcal{I})$ $:=\min\{|\mathcal{J}|\colon\mathcal{J}\subseteq\mathcal{I}\text{\ and }\bigcup\mathcal{J}=X\};$
$\operatorname{non}(\mathcal{I})$ $:=\min\{|A|\colon A\subseteq X\text{\ and }A\notin\mathcal{I}\};$
[…]
These cardinals are called additivity, covering, uniformity, and cofinality of $\mathcal{I}$, respectively. The relationship between the cardinals defined above is illustrated in Figure 1.
[…]
Figure 1. Diagram of the cardinal characteristics associated with $\mathcal{I}$. An arrow $\mathfrak{x}\rightarrow\mathfrak{y}$ means that (provably in ZFC) $\mathfrak{x}\leq\mathfrak{y}$.
Yorioka [Yor02] defined ideals $\mathcal{I}_{f}$ on $^{\omega}2$, parametrized by $f\in{}^{\uparrow\omega}\omega$, which have Borel bases and characterize $\mathcal{SN}:=\bigcap_{f\in{}^{\uparrow\omega}\omega}\mathcal{I}_{f}$. A further property of the Yorioka ideals is that $\mathcal{I}_{g}\subseteq\mathcal{I}_{f}$ whenever $f\leq^{*}g$.
These ideals led Yorioka to rediscover and significantly improve Seredyński [Ser89] result about the cofinality of $\mathcal{SN}$, and to show that no inequality between $\operatorname{cof}(\mathcal{SN})$ and $\mathfrak{c}$ can be decided in $\mathrm{ZFC}$. More research about the cofinality has been conducted by the first and second authors [Car22, CM23]. One representative result is the following, whose notation is clarified afterward.
Theorem 1.2 ([CM23, Thm. 4.6]).
$\operatorname{cof}(\mathcal{SN})\leq\operatorname{cov}\left(\left([\operatorname{supcof}(\vec{\mathcal{S}})]^{<\operatorname{minadd}(\vec{\mathcal{S}})}\right)^{\mathfrak{d}}\right).$
We discovered that the result above is not associated with the meaning of strong measure zero but of the framework given by the Yorioka ideals. We generalize this framework below, with the central notion of this paper.
Definition 1.3.
Let $X$ be a set and let $S$ be a directed preorder. We say that a sequence $\vec{\mathcal{I}}=\langle\mathcal{I}_{s}\colon s\in S\rangle$ of ideals on $X$ is a directed scheme of ideals on $X$ if $\mathcal{I}_{s^{\prime}}\subseteq\mathcal{I}_{s}$ whether $s\leq s^{\prime}$. Put $\vec{\mathcal{I}}^{*}=\bigcap_{s\in S}\mathcal{I}_{s}$, which we simply abbreviate $\mathcal{I}^{*}$ when clear from the context.
These directed schemes of ideals give rise to the study of the cardinals associated with $\mathcal{I}_{s}$. Moreover, we associate the following cardinal characteristics to $\vec{\mathcal{I}}$.
$\operatorname{minadd}(\vec{\mathcal{I}})$ $:=\min_{s\in S}\operatorname{add}(\mathcal{I}_{s}),$ $\operatorname{supadd}(\vec{\mathcal{I}})$ $:=\sup_{s\in S}\operatorname{add}(\mathcal{I}_{s}),$
$\operatorname{mincov}(\vec{\mathcal{I}})$ $:=\min_{s\in S}\operatorname{cov}(\mathcal{I}_{s}),$ $\operatorname{supcov}(\vec{\mathcal{I}})$ $:=\sup_{s\in S}\operatorname{cov}(\mathcal{I}_{s}),$
$\operatorname{minnon}(\vec{\mathcal{I}})$ $:=\min_{s\in S}\operatorname{non}(\mathcal{I}_{s}),$ $\operatorname{supnon}(\vec{\mathcal{I}})$ $:=\sup_{s\in S}\operatorname{non}(\mathcal{I}_{s}),$
$\operatorname{mincof}(\vec{\mathcal{I}})$ $:=\min_{s\in S}\operatorname{cof}(\mathcal{I}_{s}).$ $\operatorname{supcof}(\vec{\mathcal{I}})$ $:=\sup_{s\in S}\operatorname{cof}(\mathcal{I}_{s}).$
It is clear that $\operatorname{cov}(\mathcal{I}_{s})\leq\operatorname{cov}(\mathcal{I}_{s^{\prime}})$ and $\operatorname{non}(\mathcal{I}_{s^{\prime}})\leq\operatorname{non}(\mathcal{I}_{s})$ whenever $s\leq s^{\prime}$ in $S$. Then, whenever $S$ has a minimum element $s_{0}$, $\operatorname{mincov}(\vec{\mathcal{I}})=\operatorname{cov}(\mathcal{I}_{s_{0}})$ and $\operatorname{supnon}(\vec{\mathcal{I}})=\operatorname{non}(\mathcal{I}_{s_{0}})$.
For example, the Yorioka ideals form a directed scheme $\vec{\mathcal{S}}:=\langle\mathcal{I}_{f}\colon f\in{}^{\uparrow\omega}\omega\rangle$ where $^{\uparrow\omega}\omega$ is ordered by $\leq^{*}$, so $\mathcal{SN}=\vec{\mathcal{S}}^{*}$. Since $\operatorname{id}$ is a $\leq^{*}$ minimum element of $^{\uparrow\omega}\omega$, $\operatorname{mincov}(\vec{\mathcal{S}})=\operatorname{cov}(\mathcal{I}_{\operatorname{id}})$ and $\operatorname{supnon}(\vec{\mathcal{S}})=\operatorname{non}(\mathcal{I}_{\operatorname{id}})$. In this context, we also have $\operatorname{supadd}(\vec{\mathcal{S}})=\operatorname{add}(\mathcal{I}_{\operatorname{id}})$ and $\operatorname{mincof}(\vec{\mathcal{S}})=\operatorname{cof}(\mathcal{I}_{\operatorname{id}})$, see [CM19, Thm. 3.15].
Directed schemes of ideals is a crucial notion when it comes to obtaining information on the cardinals of $\mathcal{I}^{*}$ . In fact, the two first authors [Car22, CM23] used $\vec{\mathcal{S}}$ when studying the cofinality of $\mathcal{SN}$. Moreover, in [CM19], they used these ideals to investigate the rest of the cardinals associated with $\mathcal{SN}$. So one of the aims of this paper is to generalize most of the results established of the cardinals of $\mathcal{SN}$ via Yorioka ideals in the context of the directed scheme of ideals. For example, Theorem 1.2 is a consequence of our following main result, which we prove in Section 2.
Theorem A.
Let $\vec{\mathcal{I}}$ be a directed scheme of ideals on $S$. Then:
  1. (a)
    $\operatorname{minadd}(\vec{\mathcal{I}})\leq\operatorname{add}(\mathcal{I}^{*})$;
  2. (b)
    $\operatorname{supcov}(\vec{\mathcal{I}})\leq\operatorname{cov}(\mathcal{I}^{*})\leq\operatorname{cov}\left(\prod_{s\in D}\mathrm{Cv}_{\mathcal{I}_{s}}\right)$;
  3. (c)
    $\operatorname{non}(\mathcal{I}^{*})=\operatorname{minnon}(\vec{\mathcal{I}})$;
  4. (d)
    $\operatorname{cof}(\mathcal{I}^{*})\leq\mathfrak{d}\left(\prod\vec{\mathcal{I}}\right)\leq\operatorname{cov}\left(\left([\operatorname{supcof}(\vec{\mathcal{I}})]^{<\operatorname{minadd}(\vec{\mathcal{I}})}\right)^{\operatorname{cof}(S)}\right)$.
Throughout time, numerous characterizations of $\mathcal{SN}$ were discovered. One of the very interesting characterizations of $\mathcal{SN}$ is the Galvin–Mycielski–Solovay Theorem stated below, considering $^{\omega}2$ as a topological group with the standard coordinate-wise addition modulo $2$.
Theorem 1.4 ([GMS79]).
A set $X\subseteq{}^{\omega}2$ is $\mathcal{SN}$ if and only if $X+M\neq{}^{\omega}2$ for each $M\in\mathcal{M}$.
This theorem led to the rise of a certain kind of small sets, such as the meager-additive and null-additive sets, which are defined as follows.
Definition 1.5.
Let $\mathcal{I}\subseteq\mathcal{P}({}^{\omega}2)$ be an ideal.
  1. (1)
    $\mathcal{I}$ is translation invariant if $A+x\in\mathcal{I}$ for each $A\in\mathcal{I}$ and $x\in{}^{\omega}2$
  2. (2)
    A set $X\subseteq{}^{\omega}2$ is termed $\mathcal{I}$-additive if, for every $A\in\mathcal{I}$, $A+X\in\mathcal{I}$. Denote by $\mathcal{I}\mathcal{A}$ the collection of the $\mathcal{I}$-additive subsets of $^{\omega}2$. Notice that $\mathcal{I}\mathcal{A}$ is a ($\sigma$-)ideal and $\mathcal{I}\mathcal{A}\subseteq\mathcal{I}$ when $\mathcal{I}$ is a translation invariant ($\sigma$-)ideal.
  3. (3)
    The members of $\mathcal{M}\mathcal{A}$ are called meager-additive, while those in $\mathcal{N}\!\mathcal{A}$ are called null-additive.
The cardinal characteristics associated with $\mathcal{I}\mathcal{A}$ and $\mathcal{I}$ are easily related as follows.
Lemma 1.6 (e.g. [CMR25, Lem. 1.3]).
For any translation invariant ideal $\mathcal{I}$ on $^{\omega}2$:
  1. (a)
    $\operatorname{add}(\mathcal{I})\leq\operatorname{add}(\mathcal{I}\mathcal{A})$.
  2. (b)
    $\operatorname{non}(\mathcal{I}\mathcal{A})\leq\operatorname{non}(\mathcal{I})$.
  3. (c)
    $\operatorname{cov}(\mathcal{I})\leq\operatorname{cov}(\mathcal{I}\mathcal{A})$.
The ideal $\mathcal{I}\mathcal{A}$ has attracted a lot of attention when $\mathcal{I}$ is either $\mathcal{M}$ or $\mathcal{N}$. For instance, they were investigated by Bartoszyński and Judah [BJ95], Pawlikowski [Paw85], Shelah [She95], Zindulka [Zin19], and more recently by the authors [CMR25].
The ideals $\mathcal{M}\mathcal{A}$ and $\mathcal{N}\!\mathcal{A}$ are characterized below:
Theorem 1.7.
Let $X\subseteq{}^{\omega}2$.
  1. (1)
    ([She95, Thm. 13]) $X\in\mathcal{N}\!\mathcal{A}$ iff for all $I=\langle I_{n}\colon n\in\omega\rangle\in\mathbb{I}$ there is some $\varphi\in\prod_{n\in\omega}\mathcal{P}({}^{I_{n}}2)$ such that $\forall n\in\omega\colon|\varphi(n)|\leq n$ and $X\subseteq H_{\varphi}$, where
    $H_{\varphi}:=\{x\in{}^{\omega}2\colon\forall^{\infty}n\in\omega\colon x{\upharpoonright}I_{n}\in\varphi(n)\}.$
  2. (2)
    ([BJ95, Thm. 2.2]) $X\in\mathcal{M}\mathcal{A}$ iff for all $I\in\mathbb{I}$ there are $J\in\mathbb{I}$ and $y\in{}^{\omega}2$ such that
    $\forall x\in X\,\forall^{\infty}n<\omega\,\exists k<\omega\colon I_{k}\subseteq J_{n}\text{\ and\ }x{\upharpoonright}I_{k}=y{\upharpoonright}I_{k}.$
    Moreover, Shelah [She95, Thm. 18] proved that $J$ can be found coarser than $I$, i.e. every member of $J$ is the union of members of II
These characterizations led Shelah [She95] to prove that every null-additive set is meager-additive, that is $\mathcal{N}\!\mathcal{A}\subseteq\mathcal{M}\mathcal{A}$. Zindulka [Zin22], on the other hand, used combinatorial properties of meager-additive sets described by Shelah and Pawlikowski to characterize meager-additive sets in $^{\omega}2$ in a way that nicely parallels the definition of strong measure zero sets. This led him to establish the following:
Theorem 1.8.
$\mathcal{E}\mathcal{A}=\mathcal{M}\mathcal{A}$.
[…]
$\mathcal{M}\mathcal{A}=\bigcap\{\mathcal{M}_{I}\colon I\in\mathbb{I}\}.$
[…]
As in the case of the Yorioka ideals, we greatly expect that the directed scheme $\vec{\mathcal{M}}$ is useful to understand the combinatorics of $\mathcal{M}\mathcal{A}$ and its cardinal characteristics. Initially, we intended to approach the following questions stated in [CMR25].
Question 1.9.
  1. (1)
    Is $\operatorname{add}(\mathcal{M}\mathcal{A})\leq\mathfrak{b}$?
  2. (2)
    Is $\operatorname{add}(\mathcal{M}\mathcal{A})=\operatorname{non}(\mathcal{M}\mathcal{A})$?
Both questions cannot have positive answers simultaneously because $\operatorname{add}(\mathcal{M}\mathcal{A})\leq\mathfrak{b}$ implies $\operatorname{add}(\mathcal{M}\mathcal{A})=\operatorname{add}(\mathcal{M})$ (considering that $\operatorname{add}(\mathcal{M})=\min\{\mathfrak{b},\operatorname{non}(\mathcal{M}\mathcal{A})\}$, due to Pawlikowski [Paw85]), while $\mathfrak{b}<\operatorname{non}(\mathcal{M}\mathcal{A})$ is consistent with $\mathrm{ZFC}$ ([Paw85, Thm. 2.4]).
Instead of solving the previous questions, the directed scheme $\vec{\mathcal{M}}$ gives us interesting characterizations of category.
Theorem B (Theorem 3.17).
For all $I\in\mathbb{I}$, we have
$\operatorname{add}(\mathcal{M}_{I})=\operatorname{add}(\mathcal{M})$ $=\min\{\mathfrak{b},\operatorname{minnon}(\vec{\mathcal{M}})\},$
$\operatorname{cof}(\mathcal{M}_{I})=\operatorname{cof}(\mathcal{M})$ $=\max\{\mathfrak{d},\operatorname{supcov}(\vec{\mathcal{M}})\}.$
The characterization of $\operatorname{add}(\mathcal{M})$ in this theorem is already known from Pawlikowski’s $\operatorname{add}(\mathcal{M})=\min\{\mathfrak{b},\operatorname{non}(\mathcal{M}\mathcal{A})\}$ ([Paw85], and by A, $\operatorname{minnon}(\vec{\mathcal{M}})=\operatorname{non}(\mathcal{M}\mathcal{A})$).
Consequently, it is clear that $\operatorname{minadd}(\vec{\mathcal{M}})=\operatorname{supadd}(\vec{\mathcal{M}})=\operatorname{add}(\mathcal{M})$ and $\operatorname{mincof}(\vec{\mathcal{M}})=\operatorname{supcof}(\vec{\mathcal{M}})=\operatorname{cof}(\mathcal{M})$.
Thanks to the directed scheme of ideals $\vec{\mathcal{M}}$, we can estimate an upper bound of the cofinality of $\mathcal{M}\mathcal{A}$. Combining A and B results in:
Corollary 1.10.
For any $I\in\mathbb{I}$,
$\operatorname{add}(\mathcal{M})\leq\operatorname{add}(\mathcal{M}\mathcal{A})\text{\ and\ }\operatorname{cof}(\mathcal{M}\mathcal{A})\leq\operatorname{cov}(([\mathrm{cof}(\mathcal{M})]^{<\scriptstyle{\mathrm{add}(\mathcal{M})}})^{\mathfrak{d}})$
Note that the first inequality in the prior theorem is already known (1.6 (a)).
Below, we illustrate the connections between the cardinal characteristics associated with the ideals $\mathcal{M}_{I}$ and some other classical characteristics (Figure 2).
Theorem C.
For $I\in\mathbb{I}$:
  1. (1)
    If $\sum_{k<\omega}2^{-|I_{k}|}<\infty$, then $\operatorname{non}(\mathcal{M}_{I})\leq\operatorname{non}(\mathcal{E})$ and $\operatorname{cov}(\mathcal{E})\leq\operatorname{cov}(\mathcal{M}_{I})$.
  2. (2)
    If $\sum_{k<\omega}2^{-|I_{k}|}=\infty$, then $\operatorname{cov}(\mathcal{N})\leq\operatorname{non}(\mathcal{M}_{I})$ and $\operatorname{cov}(\mathcal{M}_{I})\leq\operatorname{non}(\mathcal{N})$.
$\aleph_{1}$$\operatorname{add}(\mathcal{N})$$\operatorname{cov}(\mathcal{N})$$\operatorname{non}(\mathcal{N})$$\operatorname{cof}(\mathcal{N})$${\color[rgb]{0.0,0.5,1.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,1.0}\operatorname{add}(\mathcal{M}_{I})}=\operatorname{add}(\mathcal{M})=\operatorname{add}(\mathcal{E})$$\operatorname{cov}(\mathcal{M})$$\operatorname{non}(\mathcal{M})$$\operatorname{cof}(\mathcal{E})=\operatorname{cof}(\mathcal{M})={\color[rgb]{0.0,0.5,1.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,1.0}\operatorname{cof}(\mathcal{M}_{I})}$$\mathfrak{b}$$\mathfrak{d}$$\operatorname{cov}(\mathcal{E})$$\operatorname{non}(\mathcal{E})$$\mathfrak{c}$$\operatorname{add}(\mathcal{M}\mathcal{A})$$\operatorname{non}(\mathcal{M}\mathcal{A})$$\operatorname{non}(\mathcal{M}_{I})$$\operatorname{cov}(\mathcal{M}_{I})$$\operatorname{supcov}(\vec{\mathcal{M}})$$\operatorname{cov}(\mathcal{M}\mathcal{A})$$\operatorname{cof}(\mathcal{M}\mathcal{A})$$\operatorname{cov}\left(([\operatorname{cof}(\mathcal{M})]^{<\operatorname{add}(\mathcal{M})})^{\mathfrak{d}}\right)$$2^{\mathfrak{d}}$
Figure 2. Cichoń’s diagram including the cardinal characteristics associated with our ideals, and $\operatorname{add}(\mathcal{E})=\operatorname{add}(\mathcal{M})$ and $\operatorname{cof}(\mathcal{E})=\operatorname{cof}(\mathcal{M})$ due to Bartoszyński and Shelah [BS92]. The inequality $\operatorname{non}(\mathcal{M}_{I})\leq\operatorname{non}(\mathcal{E})$ and $\operatorname{cov}(\mathcal{E})\leq\operatorname{cov}(\mathcal{M}_{I})$ holds whenever $\sum_{k<\omega}2^{-|I_{k}|}<\infty$.
B and C will be proved in Section 3. At the end of this section, we further show the connections between $\mathcal{M}_{I}$ and $\mathbf{R}_{b}$ ($b\in{}^{\omega}\omega$), a relational system that helps us in [CMR25] dissect and reformulate Barotszyński’s and Judah’s [BJ95] characterization of $\operatorname{non}(\mathcal{M}\mathcal{A})$.
In Section 4 we present consistency results. The authors proved in [CMR25] the consistency of $\operatorname{cov}(\mathcal{M}\mathcal{A})<\mathfrak{c}$, even $\operatorname{cov}(\mathcal{M}\mathcal{A})<\operatorname{non}(\mathcal{SN})$. Since $\mathcal{M}\mathcal{A}\subseteq\mathcal{SN}$, $\operatorname{cov}(\mathcal{SN})\leq\operatorname{cov}(\mathcal{M}\mathcal{A})$, but $\operatorname{cov}(\mathcal{SN})=\mathfrak{c}$ holds in Sacks model (see e.g. [CMR22]). We asked about the consistency of $\operatorname{cov}(\mathcal{N}\!\mathcal{A})<\mathfrak{c}$ in [CMR25], which we solve in this paper.
Theorem D.
It is consistent with $\mathrm{ZFC}$ that $\operatorname{cov}(\mathcal{N}\!\mathcal{A})<\mathfrak{c}$.
To prove this, we refine Shelah’s characterization of $\mathcal{N}\!\mathcal{A}$ from Theorem 1.7 (1) and show a relevant Tukey connection that gives us a suitable upper bound of $\operatorname{cov}(\mathcal{N}\!\mathcal{A})$.
A final consistency result is the following.
Theorem E.
It is consistent with $\mathrm{ZFC}$ that $\operatorname{cof}(\mathcal{M}\mathcal{A})<\mathfrak{c}$, even $\operatorname{cof}(\mathcal{M}\mathcal{A})<\operatorname{non}(\mathcal{SN})$.
The directed scheme $\vec{\mathcal{M}}$ helps us see that this inequality holds in our model of $\operatorname{cov}(\mathcal{M}\mathcal{A})<\operatorname{non}(\mathcal{SN})$ presented in [CMR25].
2. Directed schemes of ideals
[…]
  1. (1)
    A set $F\subseteq X$ is $\mathbf{R}$-bounded if $\exists\,y\in Y\ \forall\,x\in F\colon x\sqsubset y$.
  2. (2)
    A set $E\subseteq Y$ is $\mathbf{R}$-dominating if $\forall\,x\in X\ \exists\,y\in E\colon x\sqsubset y$.
We associate two cardinal characteristics with this relational system:
  • $\mathfrak{b}(\mathbf{R}):=\min\{|F|\colon F\subseteq X\text{ is }\mathbf{R}\text{-unbounded}\}$ the bounding number of $\mathbf{R}$, and
  • $\mathfrak{d}(\mathbf{R}):=\min\{|D|\colon D\subseteq Y\text{ is }\mathbf{R}\text{-dominating}\}$ the dominating number of $\mathbf{R}$.
The dual of $\mathbf{R}$ is the relational system $\mathbf{R}^{\perp}:=\langle Y,X,\sqsubset^{\perp}\rangle$ where $y\sqsubset^{\perp}x$ means $x\not\sqsubset y$. Note that $\mathfrak{b}(\mathbf{R}^{\perp})=\mathfrak{d}(\mathbf{R})$ and $\mathfrak{d}(\mathbf{R}^{\perp})=\mathfrak{b}(\mathbf{R})$.
For instance, a preorder $\langle P,\leq\rangle$ (i.e. $\leq$ is reflexive and transitive in $P$) can be viewed as the relational system $\langle P,P,\leq\rangle$, simply denoted by $P$ when $\leq$ is understood. In this context, $\mathfrak{d}(P)$ is usually denoted by $\operatorname{cof}(P)$ and $\operatorname{cf}(P)$, and called the cofinality of $P$.
The following are representative directed preorders, mostly present in this work.
Definition 2.1.
Define the relation $\leq^{*}$ on $^{\omega}\omega$ by $x\leq^{*}y$ means $\forall^{\infty}n<\omega\colon x(n)\leq y(n)$. It is clear that $\langle{}^{\omega}\omega\leq^{*}\rangle$ is a directed preorder and we use $^{\omega}\omega$ to denote its relational system. Then $\mathfrak{b}:=\mathfrak{b}({}^{\omega}\omega)$ and $\mathfrak{d}:=\mathfrak{d}({}^{\omega}\omega)$ are the well-known bounding number and dominating number, respectively.
Example 2.2.
Let $\mathcal{I}$ be an ideal on $X$ containing $X]^{<\aleph_{0}}$. As a relational system, $\mathcal{I}$ is the partial order $\langle\mathcal{I},\subseteq\rangle$, which is directed. Another relational system associated with $\mathcal{I}$ is $\mathrm{Cv}_{\mathcal{I}}:=\langle X,\mathcal{I},\in\rangle$. The cardinal characteristics associated with $\mathcal{I}$ are obtained by these relational systems:
$\mathfrak{b}(\mathcal{I})$ $=\operatorname{add}(\mathcal{I}),$ $\mathfrak{d}(\mathcal{I})$ $=\operatorname{cof}(\mathcal{I}),$
$\mathfrak{b}(\mathrm{Cv}_{\mathcal{I}})$ $=\operatorname{non}(\mathcal{I}),$ $\mathfrak{d}(\mathrm{Cv}_{\mathcal{I}})$ $=\operatorname{cov}(\mathcal{I}).$
For relational systems $\mathbf{R}=\langle X,Y,\sqsubset\rangle$ and $\mathbf{R}^{\prime}=\langle X^{\prime},Y^{\prime},\sqsubset^{\prime}\rangle$, a pair $(\Psi_{-},\Psi_{+})$ is a Tukey connection from $\mathbf{R}$ into $\mathbf{R}^{\prime}$ if $\Psi_{-}\colon X\to X^{\prime}$ and $\Psi_{+}\colon Y^{\prime}\to Y$ are functions satisfying
$\forall\,x\in X\ \forall\,y^{\prime}\in Y^{\prime}\colon\Psi_{-}(x)\sqsubset^{\prime}y^{\prime}\Rightarrow x\sqsubset\Psi_{+}(y^{\prime}).$
We say that $\mathbf{R}$ is Tukey below $\mathbf{R}^{\prime}$, denoted by $\mathbf{R}\leq_{\mathrm{T}}\mathbf{R}^{\prime}$, if there is a Tukey connection from $\mathbf{R}$ into $\mathbf{R}^{\prime}$, and we say that $\mathbf{R}$ and $\mathbf{R}^{\prime}$ are Tukey equivalent, denoted by $\mathbf{R}\cong_{\mathrm{T}}\mathbf{R}^{\prime}$, if $\mathbf{R}\leq_{\mathrm{T}}\mathbf{R}^{\prime}$ and $\mathbf{R}^{\prime}\leq_{\mathrm{T}}\mathbf{R}$. It is well-known that $\mathbf{R}\leq_{\mathrm{T}}\mathbf{R}^{\prime}$ implies $(\mathbf{R}^{\prime})^{\perp}\leq_{\mathrm{T}}\mathbf{R}^{\perp}$, $\mathfrak{b}(\mathbf{R}^{\prime})\leq\mathfrak{b}(\mathbf{R})$ and $\mathfrak{d}(\mathbf{R})\leq\mathfrak{d}(\mathbf{R}^{\prime})$. Hence, $\mathbf{R}\cong_{\mathrm{T}}\mathbf{R}^{\prime}$ implies $\mathfrak{b}(\mathbf{R}^{\prime})=\mathfrak{b}(\mathbf{R})$ and $\mathfrak{d}(\mathbf{R})=\mathfrak{d}(\mathbf{R}^{\prime})$.
Example 2.3.
  1. (1)
    Let $\mathcal{I}$ and $\mathcal{J}$ be ideals on $X$. If $\mathcal{I}\subseteq\mathcal{J}$ then $\mathrm{Cv}_{\mathcal{J}}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{I}}$.
  2. (2)
    For any ideal $\mathcal{I}$ on $X$, $\mathrm{Cv}_{\mathcal{I}}\leq_{\mathrm{T}}\mathcal{I}$ and $\mathrm{Cv}^{\perp}_{\mathcal{I}}\leq_{\mathrm{T}}\mathcal{I}$. These determine some of the inequalities in Figure 1.
  3. (3)
    If $\theta^{\prime}\leq\theta$ are cardinals and $\theta\leq|X|\leq|X^{\prime}|$, then $\mathrm{Cv}_{[X]^{<\theta}}\leq_{\mathrm{T}}\mathrm{Cv}_{[X^{\prime}]^{<\theta^{\prime}}}$ and $X]^{<\theta}\leq_{\mathrm{T}}[X^{\prime}]^{<\theta}$.
  4. (4)
    For any cardinal $\mu$, $\mathrm{Cv}_{[\mu]^{<\mu}}\leq_{\mathrm{T}}\mu\leq_{\mathrm{T}}\mathrm{Cv}^{\perp}_{[\mu]^{<\mu}}\leq_{\mathrm{T}}[\mu]^{<\mu}$. In the case when $\mu$ is regular, $\mu]^{<\mu}\leq_{\mathrm{T}}\mathrm{Cv}_{[\mu]^{<\mu}}$, so $\operatorname{add}([\mu]^{<\mu})=\operatorname{cof}([\mu]^{<\mu})=\mu$.
  5. (5)
    Whenever $S$ is a directed preorder, $S\leq_{\mathrm{T}}\mathrm{Cv}_{[\mathfrak{d}(S)]^{<\mathfrak{b}(S)}}$.
Example 2.4 ([Bla10]).
The directed preorder $\mathbb{I}=\langle\mathbb{I},\sqsubseteq\rangle$ is Tukey-equivalent with $^{\omega}\omega$. Therefore, $\mathfrak{b}(\mathbb{I})=\mathfrak{b}$ and $\mathfrak{d}(\mathbb{I})=\mathfrak{d}$.
To prove the results of this work, we employ the following operations of relational systems.
Definition 2.5.
Let $\vec{\mathbf{R}}=\langle\mathbf{R}_{i}\colon i\in K\rangle$, $\mathbf{R}_{i}=\langle X_{i},Y_{i},\sqsubset_{i}\rangle$, be a sequence of relational systems. Define the following relational systems.
  1. (1)
    $\bigvee\vec{\mathbf{R}}=\bigvee_{i\in K}\mathbf{R}_{i}:=\left\langle\bigcup_{i\in K}\{i\}\times X_{i},\prod_{i\in K}Y_{i},\sqsubset^{\vee}\right\rangle$ such that $(i,x)\sqsubset^{\vee}y$ iff $x\sqsubset_{i}y_{i}$.
  2. (2)
    $\prod\vec{\mathbf{R}}=\prod_{i\in K}\mathbf{R}_{i}:=\left\langle\prod_{i\in K}X_{i},\prod_{i\in K}Y_{i},\sqsubset^{\otimes}\right\rangle$ where $x\sqsubset^{\otimes}y$ iff $x_{i}\sqsubset_{i}y_{i}$ for all $i\in K$.
    When $\mathbf{R}_{i}=\mathbf{R}$ for all $i\in K$, we write $\mathbf{R}^{K}:=\prod\vec{\mathbf{R}}$.
  3. (3)
    When $K=\gamma$ is an ordinal, the sequential composition is
    $\vec{\mathbf{R}}^{;}:=\left\langle\prod_{\alpha<\gamma}{}^{\prod_{i<\alpha}Y_{i}}X_{\alpha},\prod_{\alpha<\gamma}Y_{\alpha},\sqsubset^{;}\right\rangle$
    where $\bar{f}=\langle f_{\alpha}\colon\alpha<\gamma\rangle\sqsubset^{;}y$ iff $f_{\alpha}(y{\upharpoonright}\alpha)\sqsubset_{\alpha}y_{\alpha}$ for all $\alpha<\gamma$.
When $K=\{0,1\}$, we denote the previous relational systems by $\mathbf{R}_{0}\vee\mathbf{R}_{1}$, $\mathbf{R}_{0}\times\mathbf{R}_{1}$ and $(\mathbf{R}_{0};\mathbf{R}_{1})$.
Blass [Bla10, Thm. 4.11] states a result about the cardinal characteristics associated with the operations above for $K=\{0,1\}$, which we generalize below.
Theorem 2.6.
Let $\vec{\mathbf{R}}=\langle\mathbf{R}_{i}\colon i\in K\rangle$ be a sequence of relational systems $\mathbf{R}_{i}=\langle X_{i},Y_{i},\sqsubset_{i}\rangle$.
  1. (a)
    $\bigvee\vec{\mathbf{R}}$ is the supremum of $\{\mathbf{R}_{i}\colon i\in K\}$ with respect to $\leq_{\mathrm{T}}$, i.e. $\mathbf{R}_{i}\leq_{\mathrm{T}}\bigvee\vec{\mathbf{R}}$ for all $i\in K$ and, whenever $\mathbf{R}^{\prime}$ is another relational system such that $\mathbf{R}_{i}\leq_{\mathrm{T}}\mathbf{R}^{\prime}$ for all $i\in K$, $\bigvee\vec{\mathbf{R}}\leq_{\mathrm{T}}\mathbf{R}^{\prime}$.
  2. (b)
    $\bigvee\vec{\mathbf{R}}\leq_{\mathrm{T}}\prod\vec{\mathbf{R}}$ and $\prod\vec{\mathbf{R}}\leq_{\mathrm{T}}\vec{\mathbf{R}}^{;}$, the latter when $K$ is an ordinal.
  3. (c)
    $\mathfrak{b}(\bigvee\vec{\mathbf{R}})=\min_{i\in K}\mathfrak{b}(\mathbf{R}_{i})$ and $\mathfrak{d}(\bigvee\vec{\mathbf{R}})=\sup_{i\in K}\mathfrak{d}(\mathbf{R}_{i})$.
  4. (d)
    $\mathfrak{b}(\prod\vec{\mathbf{R}})=\min_{i\in K}\mathfrak{b}(\mathbf{R}_{i})$ and $\sup_{i\in K}\mathfrak{d}(\mathbf{R}_{i})\leq\mathfrak{d}(\prod\vec{\mathbf{R}})\leq\prod_{i\in K}\mathfrak{d}(\mathbf{R}_{i})$. [CM25, Fact 3.9]
  5. (e)
    When $K$ is an ordinal, $\mathfrak{b}(\vec{\mathbf{R}}^{;})=\min_{i\in K}\mathfrak{b}(\mathbf{R}_{i})$ and $\mathfrak{d}(\vec{\mathbf{R}}^{;})\leq\prod_{i\in K}\mathfrak{d}(\mathbf{R}_{i})$. Equality holds when $K$ is finite.
Proof.
Without loss of generality, we can assume that $K=\gamma$ is an ordinal.
For $i\in K$, $\mathbf{R}_{i}\leq_{\mathrm{T}}\bigvee\vec{\mathbf{R}}$ is obtained by the maps $x\mapsto(i,x)$ and $y\mapsto y(i)$.
If, for each $i\in K$, $(\Psi^{i}_{-},\Psi^{i}_{+})$ is a Tukey connection for $\mathbf{R}_{i}\leq_{\mathrm{T}}\mathbf{R}^{\prime}$, then $\bigvee\vec{\mathbf{R}}\leq_{\mathrm{T}}\mathbf{R}^{\prime}$ is obtained by the maps $(i,x)\mapsto\Psi^{i}_{-}(x)$ and $y^{\prime}\mapsto\langle\Psi^{i}_{+}(y^{\prime})\colon i\in K\rangle$. This shows (a).
It is easy to show that $\mathbf{R}_{i}\leq_{\mathrm{T}}\prod\vec{\mathbf{R}}$ for all $i\in K$, so $\bigvee\vec{\mathbf{R}}\leq_{\mathrm{T}}\prod\vec{\mathbf{R}}$ by (a). On the other hand, $\prod\vec{\mathbf{R}}\leq_{\mathrm{T}}\vec{\mathbf{R}}^{;}$ via the maps $x\mapsto\langle f^{x}_{\alpha}\colon\alpha<\gamma\rangle$ such that each $f^{x}_{\alpha}$ is the constant map with value $x_{\alpha}$, and $y\mapsto y$. This shows (b).
By (a) it is clear that $\mathfrak{b}(\bigvee\vec{\mathbf{R}})\leq\mathfrak{b}(\mathbf{R}_{i})$ and $\mathfrak{d}(\mathbf{R}_{i})\leq\mathfrak{d}(\bigvee\vec{\mathbf{R}})$ for all $i\in K$, so $\mathfrak{b}(\bigvee\vec{\mathbf{R}})\leq\min_{i\in K}\mathfrak{b}(\mathbf{R}_{i})$ and $\sup_{i\in K}\mathfrak{d}(\mathbf{R}_{i})\leq\mathfrak{d}(\bigvee\vec{\mathbf{R}})$.
For each $i\in K$, pick some $\mathbf{R}_{i}$-dominating $D^{i}:=\{y^{i}_{\xi}\colon\xi<\mathfrak{d}(\mathbf{R}_{i})\}\subseteq Y_{i}$. When $\mathfrak{d}(\mathbf{R}_{i})\leq\xi<\sup_{i\in K}\mathfrak{d}(\mathbf{R}_{i})$, set $y^{i}_{\xi}:=y^{i}_{0}$. So define $y_{\xi}:=\langle y^{i}_{\xi}\colon i\in K\rangle$ for $\xi<\sup_{i\in K}\mathfrak{d}(\mathbf{R}_{i})$. Then the set $\{y_{\xi}\colon\xi<\sup_{i\in K}\mathfrak{d}(\mathbf{R}_{i})\}$ is $\bigvee\vec{\mathbf{R}}$-dominating, so $\mathfrak{d}(\bigvee\vec{\mathbf{R}})\leq\sup_{i\in K}\mathfrak{d}(\mathbf{R}_{i})$. On the other hand, it is easy to see that $\prod_{i\in K}D_{i}$ is dominating for $\prod\vec{\mathbf{R}}$ and $\vec{\mathbf{R}}^{;}$, so $\mathfrak{d}(\prod\vec{\mathbf{R}})\leq\mathfrak{d}(\vec{\mathbf{R}}^{;})\leq\prod_{i\in K}\mathfrak{d}(\mathbf{R}_{i})$.
By (b) and what we have proved so far, $\mathfrak{b}(\vec{\mathbf{R}}^{;})\leq\mathfrak{b}(\prod\vec{\mathbf{R}})\leq\mathfrak{b}(\bigvee\vec{\mathbf{R}})\leq\min_{i\in K}\mathfrak{b}(\mathbf{R}_{i})$ and $\sup_{i\in K}\mathfrak{d}(\mathbf{R}_{i})=\mathfrak{d}(\bigvee\vec{\mathbf{R}})\leq\mathfrak{d}(\prod\vec{\mathbf{R}})\leq\mathfrak{d}(\vec{\mathbf{R}}^{;})\leq\prod_{i\in K}\mathfrak{d}(\mathbf{R}_{i})$. So, to conclude (c)(d) and part of (e), it is enough to show that $\min_{i\in K}\mathfrak{b}(\mathbf{R}_{i})\leq\mathfrak{b}(\vec{\mathbf{R}}^{;})$. Let $F\subseteq\prod_{\alpha<\gamma}{}^{\prod_{i<\alpha}Y_{i}}X_{\alpha}$ be a set of size $\min_{\alpha<\gamma}\mathfrak{b}(\mathbf{R}_{\alpha})$. Find $y_{\alpha}\in Y_{\alpha}$ by recursion on $\alpha<\gamma$ such that $f_{\alpha}(y{\upharpoonright}\alpha)\sqsubset_{\alpha}y_{\alpha}$ for all $\bar{f}=\langle f_{\alpha}\colon\alpha<\gamma\rangle\in F$. This is fine because $|F|<\min_{\alpha<\gamma}\mathfrak{b}(\mathbf{R}_{\alpha})$. Therefore, $F$ is $\vec{\mathbf{R}}^{;}$-bounded by $y:=\langle y_{\alpha}\colon\alpha<\gamma\rangle$.
It remains to show the equality of the dominating numbers in (e) when $K$ is finite. Since sequential composition is associative (modulo Tukey-equivalence), it is enough to assume that $K=\{0,1\}$. So let $D\subseteq Y_{0}\times Y_{1}$ of size $\mathfrak{d}(\mathbf{R}_{0})\cdot\mathfrak{d}(\mathbf{R}_{1})$. Set $D_{0}$ as the set of $y_{0}\in Y_{0}$ such that $D_{1}^{y_{0}}:=\{y_{1}\colon(y_{0},y_{1})\in D\}$ is $\mathbf{R}_{1}$-dominating. Then $D_{0}$ is not $\mathbf{R}_{0}$-dominating, otherwise $|D_{0}|\geq\mathfrak{d}(\mathbf{R}_{0})$ and, since $|D^{y_{0}}_{1}|\geq\mathfrak{d}(\mathbf{R}_{1})$ for $y_{0}\in D_{0}$, we would get that $|D|\geq\mathfrak{d}(\mathbf{R}_{0})\cdot\mathfrak{d}(\mathbf{R}_{1})$, a contradiction.
Hence, there is some $x_{0}\in X_{0}$ such that $x_{0}\not\sqsubset_{0}y_{0}$ for all $y_{0}\in D_{0}$. Define $f_{0}\colon\{\emptyset\}\to X_{0}$ such that $f_{0}(\emptyset):=x_{0}$ and let $f_{1}\colon Y_{0}\to X_{1}$ be any map such that, whenever $y_{0}\notin D_{0}$, $f_{1}(y_{0})\not\sqsubset_{1}y_{1}$ for any $y_{1}\in D^{y_{0}}_{1}$. We then get that $\bar{f}:=(f_{0},f_{1})\not\sqsubset^{;}(y_{0},y_{1})$ for all $(y_{0},y_{1})\in D$. ∎
Remark 2.7.
The equality of the $\mathfrak{d}$-number in Theorem 2.6 (e) does not hold in general when $K$ is infinite. Consider the case when $K=\omega$ and $\mathbf{R}_{i}=\langle\omega,\neq\rangle$ for $i<\omega$. Then $\vec{\mathbf{R}}^{;}\cong_{\mathrm{T}}\prod\vec{\mathbf{R}}$ and $\mathfrak{d}(\prod\vec{\mathbf{R}})=\operatorname{cov}(\mathcal{M})$ (this is a consequence of the well-known Bartoszyński’s and Miller’s characterization of $\operatorname{cov}(\mathcal{M})$, see e.g. [CM23, Sec. 5]), but $\prod_{i<\omega}\mathfrak{d}(\mathbf{R}_{i})=\mathfrak{c}$.
To see this Tukey equivalence, notice that $\prod\vec{\mathbf{R}}=\langle{}^{\omega}\omega,\neq^{\bullet}\rangle$ where $x\neq^{\bullet}y$ means that $x(n)\neq y(n)$ for all $n<\omega$, while $\vec{\mathbf{R}}^{;}=\langle{}^{{}^{<\omega}\omega}\omega,{}^{\omega}\omega,\neq^{\bullet}\rangle$ where we interpret $\varphi\neq^{\bullet}y$ as $\varphi(y{\upharpoonright}n)\neq y(n)$ for all $n<\omega$. It is clear that $\prod\vec{\mathbf{R}}\cong_{\mathrm{T}}\langle{}^{{}^{<\omega}\omega}\omega,\neq^{\bullet}\rangle$ where $\varphi\neq^{\bullet}\psi$ is interpreted as $\varphi(s)\neq\psi(s)$ for all $s\in{}^{<\omega}\omega$.
It is enough to show that $\vec{\mathbf{R}}^{;}\leq_{\mathrm{T}}\langle{}^{{}^{<\omega}\omega}\omega,\neq^{\bullet}\rangle$. Let $F$ be the identity function on $^{{}^{<\omega}\omega}\omega$ and define $F^{\prime}\colon{}^{{}^{<\omega}\omega}\omega\to{}^{\omega}\omega$ such that $F^{\prime}(\psi):=y_{\psi}$ is defined recursively by $y_{\psi}(n):=\psi(y_{\psi}{\upharpoonright}n)$. It is clear that $(F,F^{\prime})$ is the required Tukey connection.
Another example is $\mathbf{R}_{i}:=\langle\omega,\leq\rangle$ for $i<\omega$, where $x\leq y$ means $x(n)\leq y(n)$ for all $n<\omega$, where we also get $\vec{\mathbf{R}}^{;}\cong_{\mathrm{T}}\prod\vec{\mathbf{R}}$, whose $\mathfrak{d}$-number is $\mathfrak{d}$.
The following result is a useful trick to produce Tukey connections.
Lemma 2.8 (cf. [Voj93, 4.1.6]).
For $i\in\{0,1\}$, let $\mathbf{R}_{i}=\langle X_{i},Y_{i},\sqsubset_{i}\rangle$ be a relational system, and let $\mathbf{R}=\langle S,Y,\sqsubset\rangle$ be a relational system such that $S$ is directed and satisfying that, for $s,s^{\prime}\in S$ and $y\in Y$, if $s^{\prime}\sqsubset y$ and $s\leq_{S}s^{\prime}$ then $s\sqsubset y$.
If $\mathbf{R}_{0}\leq_{\mathrm{T}}\mathbf{R}$ and $\mathbf{R}_{1}\leq_{\mathrm{T}}\mathbf{R}$, then $\mathbf{R}_{0}\times\mathbf{R}_{1}\leq_{\mathrm{T}}\mathbf{R}$.
Proof.
For $i\in\{0,1\}$, consider the Tukey connections $\Psi^{i}_{-}\colon X_{i}\to S$ and $\Psi^{i}_{+}\colon Y\to Y_{i}$ for $\mathbf{R}_{i}\leq_{\mathrm{T}}\mathbf{R}$. Define $\Psi_{-}\colon X_{0}\times X_{1}\to S$ such that $\Psi_{-}(x_{0},x_{1})$ is above $\Psi^{0}_{-}(x_{0})$ and $\Psi^{1}_{-}(x_{1})$ in $S$, and define $\Psi_{+}\colon Y\to Y_{0}\times Y_{1}$ by $\Psi_{+}(y):=(\Psi^{0}_{+}(y_{0}),\Psi^{1}_{+}(y_{1}))$. ∎
Powers of ideals can be introduced using the product of relational systems in the following ways.
Definition 2.9 ([CM23, Def. 3.14]).
Given an ideal $\mathcal{I}$ on $X$ and a set $w$, define $\mathcal{I}^{(w)}$ as the ideal on $^{w}X$ generated by the sets of the form $\prod_{i\in w}A_{i}$ with $\langle A_{i}\colon i\in w\rangle\in{}^{w}\mathcal{I}$. Denote $\operatorname{add}(\mathcal{I}^{w}):=\mathfrak{b}(\mathcal{I}^{(w)})$, $\operatorname{cof}(\mathcal{I}^{w}):=\mathfrak{d}(\mathcal{I}^{(w)})$, $\operatorname{non}(\mathcal{I}^{w}):=\mathfrak{b}(\mathrm{Cv}_{\mathcal{I}^{(w)}})$ and $\operatorname{cov}(\mathcal{I}^{w}):=\mathfrak{d}(\mathrm{Cv}_{\mathcal{I}^{(w)}})$.
Fact 2.10 ([CM23, Fact 3.15]).
Let $w$ be a set and let $\mathcal{I}$ be an ideal on $X$. Then:
  1. (a)
    $\mathcal{I}^{w}\cong_{\mathrm{T}}\mathcal{I}^{(w)}$.
  2. (b)
    $\mathrm{Cv}_{\mathcal{I}}^{w}\cong_{\mathrm{T}}\mathrm{Cv}_{\mathcal{I}^{(w)}}$.
  3. (c)
    $\operatorname{add}(\mathcal{I}^{w})=\operatorname{add}(\mathcal{I})$ and $\operatorname{non}(\mathcal{I}^{w})=\operatorname{non}(\mathcal{I})$.
  4. (d)
    $\operatorname{cov}(\mathcal{I})\leq\operatorname{cov}(\mathcal{I}^{w})\leq\operatorname{cov}(\mathcal{I})^{|w|}$ and $\operatorname{cof}(\mathcal{I})\leq\operatorname{cof}(\mathcal{I}^{w})\leq\operatorname{cof}(\mathcal{I})^{|w|}$.
We can prove A in terms of relational systems and Tukey connections. The following result is the starting point to deal with the covering and uniformity numbers.
Theorem 2.11.
Let $X$ be a set, $S$ a directed preorder and $D\subseteq S$ cofinal in $S$. Assume that $\vec{\mathcal{I}}=\langle\mathcal{I}_{s}\colon s\in S\rangle$ is a directed scheme of ideals on $X$. Then
  1. (1)
    If $s\leq s^{\prime}$, then $\mathrm{Cv}_{\mathcal{I}_{s}}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{I}_{s^{\prime}}}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{I}^{*}}$.
  2. (2)
    $\mathrm{Cv}_{\mathcal{I}^{*}}\leq_{\mathrm{T}}\prod_{s\in D}\mathrm{Cv}_{\mathcal{I}_{s}}$.
  3. (3)
    $\operatorname{supcov}(\vec{\mathcal{I}})\leq\operatorname{cov}(\mathcal{I}^{*})\leq\mathfrak{d}\left(\prod_{s\in D}\mathrm{Cv}_{\mathcal{I}_{s}}\right)$ and $\operatorname{non}(\mathcal{I}^{*})=\operatorname{minnon}(\vec{\mathcal{I}})$
Proof.
(1) follows directly by 2.3 (1), and (3) is a consequence of (1)(2) and Theorem 2.6 (d). So it remains to show (2). Define the functions $\Psi_{-}:X\to\prod_{s\in D}X$ and $\Psi_{+}:\prod_{s\in D}\mathcal{I}_{s}\to\mathcal{I}^{*}$ where $\Psi_{-}(x)$ is the constant sequence with value $x$, and $\Psi_{+}(\langle A_{s}\colon s\in D\rangle):=\bigcap_{s\in D}A_{s}$, which is in $\mathcal{I}^{*}$ because $D$ is cofinal in $S$. These maps form the desired Tukey connection. ∎
We now look at the additivity and cofinality numbers.
Theorem 2.12.
Let $\vec{\mathcal{I}}=\langle\mathcal{I}_{s}\colon s\in S\rangle$ be a directed scheme of ideals on $X$. Then, for any cofinal $D\subseteq S$,
$\mathcal{I}^{*}\leq_{\mathrm{T}}\prod_{s\in D}\mathcal{I}_{s}\leq_{\mathrm{T}}\prod_{s\in D}\mathrm{Cv}_{[\operatorname{cof}(\mathcal{I}_{s})]^{<\operatorname{add}(\mathcal{I}_{s})}}\leq_{\mathrm{T}}\mathrm{Cv}^{D}_{[\operatorname{cof}(\vec{\mathcal{I}})]^{<\operatorname{minadd}(\vec{\mathcal{I}})}}.$
As a consequence,
$\operatorname{minadd}(\vec{\mathcal{I}})\leq\operatorname{add}(\mathcal{I}^{*})$
and, whenever $D$ is a witness of $\operatorname{cof}(S)$,
$\operatorname{cof}(\mathcal{I}^{*})\leq\mathfrak{d}\left(\prod_{s\in D}\mathcal{I}_{s}\right)\leq\operatorname{cov}\left(\left([\operatorname{supcof}(\vec{\mathcal{I}})]^{<\operatorname{minadd}(\vec{\mathcal{I}})}\right)^{\operatorname{cof}(S)}\right).$
Proof.
The inequality $\mathcal{I}^{*}\leq_{\mathrm{T}}\prod_{s\in D}\mathcal{I}_{s}$ is easy to show: define $\Psi_{-}:\mathcal{I}^{*}\to\prod_{s\in D}\mathcal{I}_{s}$ such that $\Psi_{-}(A)$ is the constant sequence with value $A$, and define $\Psi_{+}:\prod_{s\in D}\mathcal{I}_{s}\to\mathcal{I}^{*}$ such that $\Psi_{+}(\langle A_{s}\colon s\in D\rangle):=\bigcap_{s\in D}A_{s}$, which is in $\mathcal{I}^{*}$ because $D$ is cofinal in $S$. Note that $(\Psi_{-},\Psi_{+})$ is the required Tukey connection.
On the other hand, using  (3) and (5) of 2.3, we obtain
$\prod_{s\in D}\mathcal{I}_{s}\leq_{\mathrm{T}}\prod_{s\in D}\mathrm{Cv}_{[\operatorname{cof}(\mathcal{I}_{s})]^{<\operatorname{add}(\mathcal{I}_{s})}}\leq_{\mathrm{T}}\mathrm{Cv}^{D}_{[\operatorname{supcof}(\vec{\mathcal{I}})]^{<\operatorname{minadd}(\vec{\mathcal{I}})}}.\qed$
A is a direct consequence of Theorem 2.11 and Theorem 2.12.
3. Directed scheme for the meager additive ideal
The purpose of this section is to introduce the directed scheme of ideals $\langle\mathcal{M}_{I}\colon I\in\mathbb{I}\rangle$, which reformulates $\mathcal{M}\mathcal{A}=\bigcap\{\mathcal{M}_{I}\colon I\in\mathbb{I}\}$, and study these ideals. In particular, we prove B and C.
We start with reviewing the following cofinal family of the meager ideal.
Definition 3.1.
Let $I\in\mathbb{I}$ and $y\in{}^{\omega}2$. Define
$B_{y,I}:=\{x\in{}^{\omega}2\colon\forall^{\infty}n\in\omega\colon x{\upharpoonright}I_{n}\neq y{\upharpoonright}I_{n}\}.$
A pair $(y,I)\in 2^{\omega}\times\mathbb{I}$ is known as a chopped real, and these are used to produce a cofinal family of meager sets. It is clear that $B_{y,I}$ is a meager subset of $^{\omega}2$ (see e.g. [Bla10]).
Theorem 3.2 (Talagrand [Tal80], see e.g. [BJS93, Prop. 13]).
For every meager set $F\subseteq{}^{\omega}2$ and $I\in\mathbb{I}$ there are $y\in{}^{\omega}2$ and $I^{\prime}\in\mathbb{I}$ such that $F\subseteq B_{y,I^{\prime}}$ and each $I^{\prime}_{n}$ is the union of finitely many $I_{k}$’s.
Inspired by ((2)) let us introduce a new ideal related to $\mathcal{M}\mathcal{A}$ as follows.
Definition 3.3.
Fix $I\in\mathbb{I}$. For $J\in\mathbb{I}$ and $y\in{}^{\omega}2$, denote
$A_{J,y}^{I}:=\{x\in{}^{\omega}2\colon\forall^{\infty}n<\omega\,\exists k<\omega\colon I_{k}\subseteq J_{n}\text{\ and\ }x{\upharpoonright}I_{k}=y{\upharpoonright}I_{k}\},$
which is an $F_{\sigma}$-meager set, and define
$\mathcal{M}_{I}:=\{X\subseteq{}^{\omega}2\colon\exists(J,y)\in\mathbb{I}\times{}^{\omega}2\colon X\subseteq A_{J,y}^{I}\}.$
Also denote $\vec{\mathcal{M}}=\langle\mathcal{M}_{I}\colon I\in\mathbb{I}\rangle$.
By Theorem 1.7 (2), it is clear that $\mathcal{M}\mathcal{A}=\bigcap_{I\in\mathbb{I}}\mathcal{M}_{I}$, but we will offer a proof in Theorem 3.13. To make sense of the previous definition and conclude that $\vec{\mathcal{M}}$ is a directed scheme of ideals, we aim to show that $\mathcal{M}_{I}$ is a $\sigma$-ideal and $\mathcal{M}_{I^{\prime}}\subseteq\mathcal{M}_{I}$ whenever $I\sqsubseteq I^{\prime}$ in $\mathbb{I}$. The former requires some work, for which we introduce the following directed preorder.
Definition 3.4.
Fix $I\in\mathbb{I}$. For $J,J^{\prime}\in\mathbb{I}$ and $y,y^{\prime}\in{}^{\omega}2$ define the relation $(J,y)\sqsubseteq^{I}(J^{\prime},y^{\prime})$ iff either $I\not\sqsubseteq J$ or, for all but finitely many $n<\omega$, there is some $\ell<\omega$ satisfying the following conditions:
  1. (i)
    $J_{\ell}$ contains some $I_{k}$, and
  2. (ii)
    for all $k<\omega$, $I_{k}\subseteq J_{\ell}$ implies $I_{k}\subseteq J^{\prime}_{n}$ and $y{\upharpoonright}I_{k}=y^{\prime}{\upharpoonright}I_{k}$.
Notice that $I\sqsubseteq J$ and $(J,y)\sqsubseteq^{I}(J^{\prime},y^{\prime})$ implies $I\sqsubseteq J^{\prime}$.
Lemma 3.5.
$\langle\mathbb{I}\times{}^{\omega}2,\sqsubset^{I}\rangle$ is a directed preorder and $\mathfrak{b}(\mathbb{I}\times{}^{\omega}2,\sqsubseteq^{I})$ is uncountable.
Proof.
It is easy to show that $\langle\mathbb{I}\times{}^{\omega}2,\sqsubseteq^{I}\rangle$ is a preorder, so it is enough to prove that $\mathfrak{b}(\mathbb{I}\times{}^{\omega}2,\sqsubseteq^{I})$ is uncountable (which implies directedness). Let $(J^{m},y^{m})\in\mathbb{I}\times{}^{\omega}2$ for $m<\omega$ and, wlog, assume that $I\sqsubseteq J^{m}$ for all $m<\omega$. Construct a $J=\langle J_{n}\colon n<\omega\rangle\in\mathbb{I}$ such that each $J_{n}$ contains one interval from $J^{m}$, say $J^{m}_{\ell_{n,m}}$, for each $m\leq n$, also demanding that $\{J^{m}_{\ell_{n,m}}\colon m\leq n\}$ is pairwise disjoint and that each member of this set contains an interval from $I$. So we can define a $y\in{}^{\omega}2$ such that $y{\restriction}J_{\ell_{m,n}}^{m}=y^{m}{\restriction}J_{\ell_{m,n}}^{m}$ for all $m\leq n<\omega$. Then $(J^{m},y^{m})\sqsubset^{I}(J,y)$ for all $m<\omega$. ∎
Main Lemma 3.6.
Let $I,J,J^{\prime}$ and $y,y^{\prime}$ as in 3.4. Then, we have that $A_{J,y}^{I}\subseteq A_{J^{\prime},y^{\prime}}^{I}$ iff $(J,y)\sqsubset^{I}(J^{\prime},y^{\prime})$.
Proof.
$\Leftarrow$" Assume that $x\in A_{J,y}^{I}$, so $\forall^{\infty}n<\omega\,\exists k<\omega\colon I_{k}\subseteq J_{n}$ and $x{\upharpoonright}I_{k}=y{\upharpoonright}I_{k}$. Since $(J,y)\sqsubset^{I}(J^{\prime},y^{\prime})$, for large enough $n$ we can choose an $\ell<\omega$ such that
  1. ($\oplus_{1}$)
    $\exists k<\omega\colon I_{k}\subseteq J_{\ell}$ and $x{\restriction}I_{k}=y{\restriction}I_{k}$, and
  2. ($\oplus_{2}$)
    $\forall k<\omega\colon I_{k}\subseteq J_{\ell}\Rightarrow I_{k}\subseteq J^{\prime}_{n}$ and $y{\restriction}I_{k}=y^{\prime}{\restriction}I_{k}$.
Observe that ($\oplus_{2}$) implies $I_{k}\subseteq J^{\prime}_{n}$ and $x{\restriction}I_{k}=y^{\prime}{\restriction}I_{k}$ for a $k$ as in ($\oplus_{1}$). Hence, $x\in A_{J^{\prime},y^{\prime}}^{I}$.
$\Rightarrow$" We prove the contrapositive. Assume $(J,y)\not\sqsubset^{I}(J^{\prime},y^{\prime})$, so $I\sqsubseteq J$, i.e. there is some $\ell_{0}<\omega$ such that $J_{\ell}$ contains some $I_{k}$ for all $\ell\geq\ell_{0}$, and there are infinitely many $n<\omega$ such that:
  1. $(\star)$
    for all $\ell\geq\ell_{0}$ there is some $k<\omega$ such that $I_{k}\subseteq J_{\ell}$ and either $I_{k}\not\subseteq J^{\prime}_{n}$ or $y{\restriction}I_{k}\neq y^{\prime}{\restriction}I_{k}$.
Let $w$ be an infinite set of $n<\omega$ satisfying $(\star)$3 and such that $|\{n\in w\colon J_{\ell}\cap J^{\prime}_{n}\neq\emptyset\}|\leq 1$ for any $\ell<\omega$. Denote by $n_{\ell}$ the unique member of this set, in case it exists.
Define $x\in{}^{\omega}2$ at each $I_{k}$ as follows: First consider the case when $I_{k}\subseteq J_{\ell}$ for some $\ell\geq\ell_{0}$ (which is unique). In the case $n_{\ell}$ exists, there is some $k_{\ell}<\omega$ such that $I_{k_{\ell}}\subseteq J_{\ell}$ and either $I_{k_{\ell}}\not\subseteq J^{\prime}_{n_{\ell}}$ or $y{\restriction}I_{k_{\ell}}\neq y^{\prime}{\restriction}I_{k_{\ell}}$. Define $x{\restriction}I_{k}:=y{\restriction}I_{k}$ when $k=k_{\ell}$, otherwise $x{\restriction}I_{k}$ can be anything different from $y^{\prime}{\restriction}I_{k}$; in the case that $n_{\ell}$ does not exist, set $x{\restriction}I_{k}:=y{\restriction}I_{k}$. On the other hand, in the case that $I_{k}$ is not contained in any $J_{\ell}$ for $\ell\geq\ell_{0}$, let $x{\restriction}I_{k}$ be anything different from $y^{\prime}{\restriction}I_{k}$.
Firstly, it is clear that $x\in A_{J,y}^{I}$. On the other hand, $\forall n\in w\,\forall k<\omega\colon I_{k}\subseteq J^{\prime}_{n}\Rightarrow x{\restriction}I_{k}\neq y^{\prime}{\restriction}I_{k}$, i.e, $x\not\in A_{J^{\prime},y^{\prime}}^{I}$. Indeed, for $n\in w$ and $k<\omega$ such that $I_{k}\subseteq J^{\prime}_{n}$, in case $I_{k}\subseteq J_{\ell}$ for some $\ell\geq\ell_{0}$, we have $n=n_{\ell}$ and, if $k=k_{\ell}$, then $x{\restriction}I_{k_{\ell}}=y{\restriction}I_{k_{\ell}}\neq y^{\prime}{\restriction}I_{k_{\ell}}$, else $x{\restriction}I_{k}\neq y^{\prime}{\restriction}I_{k}$ by the definition of $x{\restriction}I_{k}$; otherwise, $I_{k}$ is not contained in any $J_{\ell}$ for $\ell\geq\ell_{0}$, in which case $x{\restriction}I_{k}\neq y^{\prime}{\restriction}I_{k}$. ∎
As a direct consequence of 3.6, not only do we have that any $\mathcal{M}_{I}$ is a $\sigma$-ideal but also that $\mathcal{M}_{I}$ is Tukey-equivalent to a clean directed preorder.
Theorem 3.7.
For $\in\mathbb{I}, I\mathcal{M}_{I}$ is a $\sigma$-ideal and $\mathcal{M}_{I}\cong_{\mathrm{T}}\langle\mathbb{I}\times{}^{\omega}2,\sqsubset^{I}\rangle$.
We must show, of course, that $^{\omega}2\notin\mathcal{M}_{I}$, but this is immediate from $\mathcal{M}_{I}\subseteq\mathcal{M}$, which we prove in 3.12.
It is clear that the set $\{(J,y)\in\mathbb{I}\times{}^{\omega}2\colon J\text{ is coarser that }I\}$ is cofinal in $\langle\mathbb{I}\times{}^{\omega}2,\sqsubset^{I}\rangle$. As a consequence:
Corollary 3.8.
We can rewrite
[…]
We point out some simple facts below, which let us conclude that $\vec{\mathcal{M}}$ is a directed scheme of ideals.
Fact 3.9.
Let $I,I^{\prime},J,J^{\prime}\in\mathbb{I}$ and $x,y,y^{\prime}\in{}^{\omega}2$. Then:
  1. (a)
    $x\in A_{J,y}^{I}$ iff $(I,x)\sqsubset^{I}(J,y)$.
  2. (b)
    If $I\sqsubseteq I^{\prime}$ then $A^{I^{\prime}}_{J,y}\subseteq A^{I}_{J,y}$. In particular, $\mathcal{M}_{I^{\prime}}\subseteq\mathcal{M}_{I}$.
We ask:
Question 3.10.
Does $I\sqsubseteq I^{\prime}$ imply $\mathcal{M}_{I^{\prime}}\leq_{\mathrm{T}}\mathcal{M}_{I}$?
Although the answer to the previous question is unknown to us, we have a stronger relation on $\mathbb{I}$ from where $\mathcal{M}_{I^{\prime}}\leq_{\mathrm{T}}\mathcal{M}_{I}$ holds. Namely, for $^{\prime}\in\mathbb{I}, I+II\sqsubseteq^{+}I^{\prime}$ means that, for all but finitely many $n<\omega$, $I^{\prime}_{n}$ is the union of some $I_{k}$’s. Notice that $\langle I^{+},\sqsubseteq^{+}\rangle$ is a preorder but not directed, even more, $\operatorname{cof}(\mathbb{I},\sqsubseteq^{+})=\mathfrak{c}$ (see [CMŠ26, Prop. 3.7]).
Fact 3.11.
For $I,I^{\prime}\in\mathbb{I}$, if $I\sqsubseteq^{+}I^{\prime}$ then $(J,y)\sqsubset^{I}(J^{\prime},y^{\prime})$ implies $(J,y)\sqsubset^{I^{\prime}}(J^{\prime},y^{\prime})$, In particular, $\langle\mathbb{I}\times{}^{\omega}2,\sqsubset^{I^{\prime}}\rangle\leq_{\mathrm{T}}\langle\mathbb{I}\times{}^{\omega}2,\sqsubset^{I}\rangle$ and $\mathcal{M}_{I^{\prime}}\leq_{\mathrm{T}}\mathcal{M}_{I}$.
Recall that $I^{1}$ denotes the partition of $\omega$ into singletons.
Lemma 3.12.
  1. (a)
    $\mathcal{M}_{I^{1}}=\mathcal{M}$.
  2. (b)
    For all $\in\mathbb{I}, I\mathcal{M}_{I}\subseteq\mathcal{M}$ and $\mathcal{M}_{I}\leq_{\mathrm{T}}\mathcal{M}$.
Proof.
We just prove (a) because (b) is consequence of (a), 3.9 and 3.11. For $x\in{}^{\omega}2$, $x\in A_{J,y}^{I^{1}}$ iff $\forall^{\infty}n\,\exists k\in J_{n}\colon x(k)=y(k)$, which is equivalent to $\forall^{\infty}n\colon x{\restriction}J_{n}\neq y^{\prime}{\restriction}J_{n}$ where $y^{\prime}:=y+1\mod{2}$. This actually shows $A_{J,y}^{I^{1}}=B_{y^{\prime},J}$ (see 3.1). Therefore, $\mathcal{M}_{I^{1}}=\mathcal{M}$. ∎
The previous representation of $\mathcal{M}$ can be used to prove:
Theorem 3.13 ([BJ95]).
$\mathcal{M}\mathcal{A}=\bigcap_{I\in\mathbb{I}}\mathcal{M}_{I}$.
Proof.
For $X\subseteq{}^{\omega}2$,
$X\in\mathcal{M}\mathcal{A}$ $\text{\ iff\ }\forall A\in\mathcal{M}\colon X+A\in\mathcal{M}$
$\text{\ iff\ }\forall(I,y)\in\mathbb{I}\times{}^{\omega}2\,\exists(J,y^{\prime})\in\mathbb{I}\times{}^{\omega}2\colon X+A_{I,y}^{I^{1}}=\underbrace{\bigcup_{x\in X}x+A_{I,y}^{I^{1}}}_{\bigcup_{x\in X}A_{I,x+y}^{I^{1}}}\subseteq A_{J,y^{\prime}}^{I^{1}}$
$\text{\ iff\ }\forall(I,y)\in\mathbb{I}\times{}^{\omega}2\,\exists(J,y^{\prime})\in\mathbb{I}\times{}^{\omega}2\,\forall x\in X\colon A_{I,x+y}^{I^{1}}\subseteq A_{J,y^{\prime}}^{I^{1}}$
$\text{\ iff\ }\forall I\in\mathbb{I}\,\exists(J,y)\in\mathbb{I}\times{}^{\omega}2\,\forall x\in X\colon A_{I,x}^{I^{1}}\subseteq A_{J,y}^{I^{1}}\text{ (using a translation)}$
$\text{\ iff\ }\forall I\in\mathbb{I}\,\exists(J,y)\in\mathbb{I}\times{}^{\omega}2\,\forall x\in X\,\forall^{\infty}n\in\omega\,\exists\ell<\omega\colon I_{\ell}\subseteq J_{n}\,\wedge\,x{\restriction}I_{\ell}=y{\restriction}I_{\ell}$
$\text{\ iff\ }\forall I\in\mathbb{I}\,\exists(J,y)\in\mathbb{I}\times{}^{\omega}2\colon X\subseteq A_{J,y}^{I}$
$\text{\ iff\ }\forall I\in\mathbb{I}\colon X\in\mathcal{M}_{I}$
$\text{\ iff\ }X\in\bigcap_{I\in\mathbb{I}}\mathcal{M}_{I}.\qed$
We now look at the cardinal characteristics associated with $\vec{\mathcal{M}}$ and its influence on those with $\mathcal{M}\mathcal{A}$ as well. Since $\vec{\mathcal{M}}$ is a directed scheme of ideals and $\mathcal{M}\mathcal{A}=\vec{\mathcal{M}}^{*}$, by Theorem 2.11 and 2.12 we immediately obtain:
Lemma 3.14.
Let $I,J\in\mathbb{I}$ and $D\subseteq\mathbb{I}$ cofinal.
  1. (a)
    If $I\sqsubseteq J$ then $\mathrm{Cv}_{\mathcal{M}}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}_{I}}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}_{J}}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}\mathcal{A}}\leq_{\mathrm{T}}\prod_{I\in D}\mathrm{Cv}_{\mathcal{M}_{I}}$. In particular, $\operatorname{cov}(\mathcal{M})\leq\operatorname{cov}(\mathcal{M}_{I})\leq\operatorname{cov}(\mathcal{M}_{J})\leq\operatorname{supcov}(\vec{\mathcal{M}})\leq\operatorname{cov}(\mathcal{M}\mathcal{A})\leq\mathfrak{d}\left(\prod_{I\in D}\mathrm{Cv}_{\mathcal{M}_{I}}\right)$ and $\operatorname{non}(\mathcal{M}\mathcal{A})=\operatorname{minnon}(\vec{\mathcal{M}})\leq\operatorname{non}(\mathcal{M}_{I})\leq\operatorname{non}(\mathcal{M}_{J})\leq\operatorname{non}(\mathcal{M})$.
  2. (b)
    $\mathcal{M}\mathcal{A}\leq_{\mathrm{T}}\prod_{I\in D}\mathcal{M}_{I}$.
It is also clear that $\operatorname{mincov}(\vec{\mathcal{M}})=\operatorname{cov}(\mathcal{M})$ and $\operatorname{supnon}(\vec{\mathcal{M}})=\operatorname{non}(\mathcal{M})$. Although 3.14 (b) gives us some information about the additivity and cofinality numbers, we are going to prove much more. Towards this direction, we first show a connection between $^{\omega}\omega$ and $\mathcal{M}_{I}$.
Lemma 3.15.
$^{\omega}\omega\leq_{\mathrm{T}}\mathcal{M}_{I}$.
Proof.
It suffices to show that $\mathbb{I}\leq_{\mathrm{T}}\langle\mathbb{I}\times{}^{\omega}2,\sqsubset^{I}\rangle$ because $\mathbb{I}\cong_{\mathrm{T}}{}^{\omega}\omega$ (see 2.4). Define the functions $\Psi_{-}\colon\mathbb{I}\to\mathbb{I}\times{}^{\omega}2$ and $\Psi_{+}\colon\mathbb{I}\times{}^{\omega}2\to\mathbb{I}$ as follows. For $J\in\mathbb{I}$, pick some $J^{+}\in\mathbb{I}$, coarser that $I$, such that $J\sqsubseteq J^{+}$.
So define
$\Psi_{-}(J)$ $:=(J^{+},\bar{0}),$
$\Psi_{+}(J^{\prime},y^{\prime})$ $:=J^{\prime},$
where $\bar{0}$ is the constant sequence with value 0.
The pair $(\Psi_{-},\Psi_{+})$ is the desired Tukey connection: Assume $(J^{+},\bar{0})\sqsubset^{I}(J^{\prime},y^{\prime})$. Since $J^{+}$ is coarser than $I$, we obtain $J^{+}\sqsubseteq J^{\prime}$, so $J\sqsubseteq J^{\prime}$. ∎
Another relevant result is the following.
Lemma 3.16.
For any $I,I^{\prime}\in\mathbb{I}$, $\mathrm{Cv}_{\mathcal{M}_{I}}\leq_{\mathrm{T}}\mathcal{M}_{I^{\prime}}$.
Proof.
Find $I^{\prime\prime}\in\mathbb{I}$ coarser than $I^{\prime}$ such that $I\sqsubseteq I^{\prime\prime}$. Using 3.11 and 3.14 (a), we get $\mathrm{Cv}_{\mathcal{M}_{I}}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}_{I^{\prime\prime}}}\leq_{\mathrm{T}}\mathcal{M}_{I^{\prime\prime}}\leq_{\mathrm{T}}\mathcal{M}_{I^{\prime}}$
We are ready to prove B. Since $\operatorname{minnon}(\vec{\mathcal{M}})=\operatorname{non}(\mathcal{M}\mathcal{A})$, it is already known from [Paw85, Lem. 2.3] that $\operatorname{add}(\mathcal{M})=\min\{\mathfrak{b},\operatorname{minnon}(\vec{\mathcal{M}})\}$.
Theorem 3.17.
For all $I\in\mathbb{I}$,
$\operatorname{add}(\mathcal{M}_{I})$ == $\operatorname{add}(\mathcal{M})$ == $\min\{\mathfrak{b},\operatorname{minnon}(\vec{\mathcal{M}})\}$,
$\operatorname{cof}(\mathcal{M}_{I})$ == $\operatorname{cof}(\mathcal{M})$ == $\max\{\mathfrak{d},\operatorname{supcov}(\vec{\mathcal{M}})\}$.
In terms of Tukey connections, $^{\omega}\omega\times\bigvee_{J\in\mathbb{I}}\mathrm{Cv}_{\mathcal{M}_{J}}\leq_{\mathrm{T}}\mathcal{M}_{I}\leq_{\mathrm{T}}\mathcal{M}\leq_{\mathrm{T}}\left(\mathbb{I};\bigvee_{J\in\mathbb{I}}\mathrm{Cv}_{\mathcal{M}_{J}}\right)$.
Proof.
Thanks to Theorem 2.6, it is enough to prove the Tukey connections. Since we already have $\mathcal{M}_{I}\leq_{\mathrm{T}}\mathcal{M}$ by 3.12, the following inequalities remain.
$^{\omega}\omega\times\bigvee_{J\in\mathbb{I}}\mathrm{Cv}_{\mathcal{M}_{J}}\leq_{\mathrm{T}}\mathcal{M}_{I}$: By 3.16 and Theorem 2.6, $\bigvee_{J\in\mathbb{I}}\mathrm{Cv}_{\mathcal{M}_{J}}\leq_{\mathrm{T}}\mathcal{M}_{I}$. On the other hand, $^{\omega}\omega\leq_{\mathrm{T}}\mathcal{M}_{I}$ by 3.15, so we can use 2.8 to conclude the desired inequality.
[…]
As a consequence, we obtain the following inequalities (although $\operatorname{add}(\mathcal{M})\leq\operatorname{add}(\mathcal{M}\mathcal{A})$ is easy to prove directly).
Corollary 3.18.
For any dominating $D\subseteq\mathbb{I}$,
$\mathcal{M}\mathcal{A}\leq_{\mathrm{T}}\prod_{I\in D}\mathcal{M}_{I}\leq_{\mathrm{T}}\mathcal{M}^{D}.$
In particular, $\mathcal{M}\mathcal{A}\leq_{\mathrm{T}}\mathcal{M}^{\mathfrak{d}}\leq_{\mathrm{T}}\mathrm{Cv}_{[\operatorname{cof}(\mathcal{M})]^{<\operatorname{add}(\mathcal{M})}}^{\mathfrak{d}}$, $\operatorname{add}(\mathcal{M})\leq\operatorname{add}(\mathcal{M}\mathcal{A})$ and $\operatorname{cof}(\mathcal{M}\mathcal{A})\leq\mathfrak{d}\left(\mathcal{M}^{\mathfrak{d}}\right)\leq\operatorname{cov}\left(([\operatorname{cof}(\mathcal{M})]^{<\operatorname{add}(\mathcal{M})})^{\mathfrak{d}}\right)$.
Next, we show how $\mathcal{M}_{I}$ is related with the measure zero ideal. From now on, $\operatorname{Lb}$ denotes the Lebesgue measure on $^{\omega}2$.
Lemma 3.19.
For $I\in\mathbb{I}$, the following statements are equivalent.
  1. (i)
    $\mathcal{M}_{I}\subseteq\mathcal{N}$.
  2. (ii)
    $\mathcal{M}_{I}\subseteq\mathcal{E}$.
  3. (iii)
    $\sum_{k<\omega}2^{-|I_{k}|}<\infty$.
Proof.
[…]
$A_{J,y}^{I}=\bigcup_{m<\omega}\bigcap_{n\geq m}\bigcup_{\begin{subarray}{c}k<\omega\\ I_{k}\subseteq J_{n}\end{subarray}}[y{\restriction}I_{k}].$
$\Leftarrow$" Assume that $\sum_{k<\omega}2^{-|I_{k}|}<\infty$. Let $X\in\mathcal{M}_{I}$, so there are $J\in\mathbb{I}$ and $y\in{}^{\omega}2$ such that $X\subseteq A_{J,y}^{I}$. To see that $X\in\mathcal{N}$, it is sufficient to show that
$\operatorname{Lb}\left(\bigcup_{m<\omega}\bigcap_{n\geq m}\bigcup_{\begin{subarray}{c}k<\omega\\ I_{k}\subseteq J_{n}\end{subarray}}[y{\restriction}I_{k}]\right)=0.$
Observe that
$\operatorname{Lb}\left(\bigcup_{\begin{subarray}{c}k<\omega\\ I_{k}\subseteq J_{n}\end{subarray}}[y{\restriction}I_{k}]\right)\leq\sum_{\begin{subarray}{c}k<\omega\\ I_{k}\subseteq J_{n}\end{subarray}}2^{-|I_{k}|}.$
Since $\sum_{k<\omega}2^{-|I_{k}|}<\infty$, $\bigcap_{n\geq m}\bigcup_{\begin{subarray}{c}k<\omega\\ I_{k}\subseteq J_{n}\end{subarray}}[y{\restriction}I_{k}]$ has measure zero. Therefore, $A_{J,y}^{I}\in\mathcal{N}$.
$\Rightarrow$" Assume that $\sum_{k<\omega}2^{-|I_{k}|}=\infty$. Let $J\in\mathbb{I}$ be coarser than $I$ such that
$\sum_{\begin{subarray}{c}k<\omega\\ I_{k}\subseteq J_{n}\end{subarray}}2^{-|I_{k}|}\geq n.$
We prove that $A_{J,y}^{I}\notin\mathcal{N}$, in fact $\operatorname{Lb}({}^{\omega}2\smallsetminus A_{J,y}^{I})=0$. Notice that
$^{\omega}2\smallsetminus A_{J,y}^{I}=\bigcap_{m<\omega}\bigcup_{n\geq m}\bigcap_{\begin{subarray}{c}k<\omega\\ I_{k}\subseteq J_{n}\end{subarray}}({}^{\omega}2\smallsetminus[y{\restriction}I_{k}]).$
To show $\operatorname{Lb}({}^{\omega}2\smallsetminus A_{J,y}^{I})=0$, it is enough to see that
$\lim_{m\to\infty}\operatorname{Lb}\left(\bigcup_{n\geq m}\bigcap_{\begin{subarray}{c}k<\omega\\ I_{k}\subseteq J_{n}\end{subarray}}{}^{\omega}2\smallsetminus[y{\restriction}I_{k}]\right)=0.$
Indeed,
$\operatorname{Lb}\left(\bigcap_{\begin{subarray}{c}k<\omega\\ I_{k}\subseteq J_{n}\end{subarray}}{}^{\omega}2\smallsetminus[y{\restriction}I_{k}]\right)$ $=\prod_{\begin{subarray}{c}k<\omega\\ I_{k}\subseteq J_{n}\end{subarray}}(1-2^{-|I_{n}|})$
$\leq e^{-\sum_{\begin{subarray}{c}k<\omega\\ I_{k}\subseteq J_{n}\end{subarray}}2^{-|I_{k}|}}$
$\leq e^{-n}$
Since $\sum_{n<\omega}e^{-n}<\infty$, we conclude the desired limit. ∎
The previous proof allows to say a bit more in the case $\sum_{k<\omega}2^{-|I_{k}|}=\infty$.
Lemma 3.20.
If $\sum_{k<\omega}2^{-|I_{k}|}=\infty$ then $\mathrm{Cv}_{\mathcal{N}}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}_{I}}^{\perp}$. In particular, $\operatorname{cov}(\mathcal{N})\leq\operatorname{non}(\mathcal{M}_{I})$ and $\operatorname{cov}(\mathcal{M}_{I})\leq\operatorname{non}(\mathcal{N})$.
Proof.
In the proof of 3.19, we produced a set $X\in\mathcal{M}_{I}$ such that $^{\omega}2\smallsetminus X\in\mathcal{N}$. Since both ideals $\mathcal{M}_{I}$ and $\mathcal{N}$ are translation invariant, we obtain $\mathrm{Cv}_{\mathcal{N}}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}_{I}}^{\perp}$ as in Rothberger’s proof of $\mathrm{Cv}_{\mathcal{N}}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}}^{\perp}$, namely, using the maps $x\mapsto x+X$ and $y\mapsto y+({}^{\omega}2\smallsetminus X)$. ∎
Using 3.19, and 3.20, C is concluded.
We finish this section with connections between our ideals and the relational systems introduced in [CMR25] to study $\mathrm{Cv}_{\mathcal{M}\mathcal{A}}$.
Definition 3.21 ([CMR25, Def. 2.7]).
Fix $b\in{}^{\omega}\omega$.
  1. (1)
    For $I\in\mathbb{I}$ and $f,h\in{}^{\omega}\omega$, define
    $f\sqsubset^{\bullet}(I,h)\textrm{\ iff\ }\forall^{\infty}n\in\omega\,\exists k\in I_{n}\colon f(k)=h(k).$
  2. (2)
    Define the relational system $\mathbf{R}_{b}:=\langle\prod b,\mathbb{I}\times\prod b,\sqsubset^{\bullet}\rangle$, where $\prod b:=\prod_{n<\omega}b(n)$.
In the context of $\mathbf{R}_{b}$, we will always consider that $b(n)>0$ for all $n$, even if we just write $b\in{}^{\omega}\omega$”.
Notice that, for fixed $(I,h)\in\mathbb{I}\times\prod b$, $\{f\in\prod b\colon f\sqsubset^{\bullet}(I,h)\}$ is meager whenever $b\geq^{*}2$, so $\mathrm{Cv}_{\mathcal{M}}\leq_{\mathrm{T}}\mathbf{R}_{b}$, which implies $\mathfrak{b}(\mathbf{R}_{b})\leq\operatorname{non}(\mathcal{M})$ and $\operatorname{cov}(\mathcal{M})\leq\mathfrak{d}(\mathbf{R}_{b})$. [^1] In [Car25], these are denoted by $\mathfrak{b}_{b}^{\mathsf{eq}}$ and $\mathfrak{d}_{b}^{\mathsf{eq}}$, respectively. On the other hand, if $b\ngeq^{*}2$ then we can find some $(I,h)\in\mathbb{I}\times\prod b$ such that $f\sqsubset^{\bullet}(I,h)$ for all $f\in\prod b$, so $\mathfrak{d}(\mathbf{R}_{b})=1$ and $\mathfrak{b}(\mathbf{R}_{b})$ is undefined.
Fact 3.22.
For $b\in\omega^{\omega}$, $\mathbf{R}_{b}\cong_{\mathrm{T}}\langle\prod b,\mathbb{I}\times{}^{\omega}\omega,\sqsubset^{\bullet}\rangle$. As a consequence, if $b^{\prime}\in\omega^{\omega}$ and $b\leq^{*}b^{\prime}$, then $\mathbf{R}_{b}\leq_{\mathrm{T}}\mathbf{R}_{b^{\prime}}$. In particular, $\mathfrak{b}(\mathbf{R}_{b^{\prime}})\leq\mathfrak{b}(\mathbf{R}_{b})$ and $\mathfrak{d}(\mathbf{R}_{b})\leq\mathfrak{d}(\mathbf{R}_{b^{\prime}})$.
In [CMR25] we have proved that $\mathbf{R}_{b}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}\mathcal{A}}\leq_{\mathrm{T}}\prod_{b\in D}\mathbf{R}_{b}$ for any $b\in{}^{\omega}\omega$ and any dominating $D\subseteq{}^{\omega}\omega$, and also that $\mathfrak{d}(\mathbf{R}_{b})\leq\operatorname{cof}(\mathcal{M})$. These facts can be deduced from the following result that connects $\mathbf{R}_{b}$ and $\mathrm{Cv}_{\mathcal{M}_{I}}$.
Lemma 3.23.
For $I\in\mathbb{I}$, if $\forall^{\infty}k\colon b(k)=2^{|I_{k}|}$ then $\mathbf{R}_{b}\cong_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}_{I}}$.
Proof.
Assume wlog that, for all $k\in\omega$, $b(k)=2^{|I_{k}|}$, so $\prod b=2^{I}:=\prod_{k<\omega}{}^{I_{k}}2$. Define
$f\colon 2^{I}$ $\longrightarrow{}^{\omega}2$
$\eta$ $\longmapsto\bigcup_{n<\omega}\eta(n)$
Note that $f$ is a bijection.
Let $\mathbb{I}_{I}$ be the set of $J\in\mathbb{I}$ coarser than $I$. Define the maps
$g\colon\mathbb{I}_{I}$ $\longrightarrow\mathbb{I}\text{ and}$
$h\colon\mathbb{I}\times 2^{I}$ $\longrightarrow\mathbb{I}_{I}\times{}^{\omega}2$
by the assignments
$J$ $\longmapsto\langle\{k<\omega\colon I_{k}\subseteq J_{n}\}\colon n<\omega\rangle\text{ and}$
$(J,y)$ $\longmapsto\big(g^{-1}(J),f(y)\big),$
respectively. These maps are bijections as well.
To define $\Psi_{+}\colon\mathcal{M}_{I}\to\mathbb{I}\times 2^{I}$, for $X\in\mathcal{M}_{I}$, pick an $(J,y)\in\mathbb{I}_{I}\times{}^{\omega}2$ such that $X\subseteq A_{J,y}^{I}$ and put $\Psi_{+}(X):=h^{-1}(J,y)$. Notice that $f(\eta)\in A_{J,y}^{I}$ iff $\forall^{\infty}n\,\exists k\colon I_{k}\subseteq J_{n}$ and $\eta(k)=y{\restriction}I_{k}$, which is equivalent to $\forall^{\infty}n\,\exists k\in J_{n}^{\prime}\colon\eta(k)=y{\restriction}I_{k}$ where $J^{\prime}=g(J)$, i.e. $\eta\sqsubset^{\bullet}\Psi_{+}(X)$. Then $(f,\Psi_{+})$ witnesses $\mathbf{R}_{b}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}_{I}}$.
To prove $\mathrm{Cv}_{\mathcal{M}_{I}}\leq_{\mathrm{T}}\mathbf{R}_{b}$, define
$\Psi^{\prime}_{+}\colon\mathbb{I}\times 2^{I}$ $\rightarrow\mathcal{M}_{I}$
$(J^{\prime},\eta)$ $\mapsto A_{(g^{-1}(J^{\prime}),f(\eta))}^{I}$
It is clear that $(f^{-1},\Psi^{\prime}_{+})$ witnesses $\mathrm{Cv}_{\mathcal{M}_{I}}\leq_{\mathrm{T}}\mathbf{R}_{b}$. Indeed, $f^{-1}(x)\sqsubset^{\bullet}(J^{\prime},\eta)$ iff $\forall^{\infty}n\,\exists k\in J^{\prime}_{n}\colon x{\restriction}I_{k}=\eta(k)$, which is equivalent to $\forall^{\infty}n\,\exists k\colon I_{k}\subseteq J_{n}$ and $x{\restriction}I_{k}=\eta(k)$ , i.e. $x\in A_{(g^{-1}(J^{\prime}),f(\eta))}^{I}$. ∎
Because of the above result, we obtain:
Corollary 3.24.
  1. (1)
    If $\forall^{\infty}k\colon\log_{2}b(k)\leq|I_{k}|$, then $\mathbf{R}_{b}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}_{I}}$.
  2. (2)
    If $\forall^{\infty}k\colon|I_{k}|\leq\log_{2}b(k)$, then $\mathrm{Cv}_{\mathcal{M}_{I}}\leq_{\mathrm{T}}\mathbf{R}_{b}$.
Corollary 3.25.
$\min_{b\in{}^{\omega}\omega}\mathfrak{b}(\mathbf{R}_{b})=\operatorname{minnon}(\vec{\mathcal{M}})$ and $\sup_{b\in{}^{\omega}\omega}\mathfrak{d}(\mathbf{R}_{b})=\operatorname{supcov}(\vec{\mathcal{M}})$. [^2] The equality $\min_{b\in{}^{\omega}\omega}\mathfrak{b}(\mathbf{R}_{b})=\operatorname{non}(\mathcal{M}\mathcal{A})$ is originally due to Bartoszyński and Judah [BJ95].
4. Consistency results
In this section, we prove our consistency results, i.e. D-E. We start reviewing some notation.
Definition 4.1.
Given a sequence of non-empty sets $b=\langle b(n)\colon n\in\omega\rangle$ and $h\colon\omega\to\omega$, define
$\prod b$ $:=\prod_{n\in\omega}b(n),\textrm{\ and}$
$\mathcal{S}(b,h)$ $:=\prod_{n\in\omega}[b(n)]^{\leq h(n)}.$
For two functions $x\in\prod b$ and $\varphi\in\mathcal{S}(b,h)$ write
[…]
Let $\mathrm{Lc}(b,h):=\langle\prod b,\mathcal{S}(b,h),\in^{*}\rangle$, which is a relational system. Denote $\mathfrak{b}^{\mathrm{Lc}}_{b,h}=\mathfrak{b}(\mathrm{Lc}(b,h))$ and $\mathfrak{d}^{\mathrm{Lc}}_{b,h}=\mathfrak{d}(\mathrm{Lc}(b,h))$. [^3] In [KM22, CKM24], these are denoted by $\mathfrak{v}_{b,h}^{\forall}$ and $\mathfrak{c}_{b,h}^{\forall}$, respectively. Also define
$\mathrm{minLc}:=\min\{\mathfrak{b}^{\mathrm{Lc}}_{b,\operatorname{id}}\colon b\in{}^{\omega}\omega\}\text{\ and\ }\mathrm{supLc}:=\sup\{\mathfrak{d}^{\mathrm{Lc}}_{b,\operatorname{id}}\colon b\in{}^{\omega}\omega\}.$
Recall the following characterization. Pawlikowski [Paw85] characterized $\operatorname{add}(\mathcal{N})$, while $\operatorname{cof}(\mathcal{N})$ is due to the first and second author [CM19, Lem. 3.11].
Theorem 4.2.
$\operatorname{add}(\mathcal{N})=\min\{\mathfrak{b},\mathrm{minLc}\}$ and $\operatorname{cof}(\mathcal{N})=\max\{\mathfrak{d},\mathrm{supLc}\}$.
Below, we strengthen the characterization of $\mathcal{N}\!\mathcal{A}$ from Theorem 1.7 (1).
Theorem 4.3.
Let $X\subseteq{}^{\omega}2$ and $D\subseteq\mathbb{I}$ a dominating family. Then $X\in\mathcal{N}\!\mathcal{A}$ iff, for all $I\in D$, there is some $\varphi\in\prod_{n\in\omega}\mathcal{P}({}^{I_{n}}2)$ such that $\forall n\in\omega\colon|\varphi(n)|\leq n$ and $X\subseteq H_{\varphi}$.
Proof.
Under Theorem 1.7 (1), we only need to prove $\Leftarrow$”, that is, the statement about $D$ implies the same statement but for $\mathbb{I}$ in the place of $D$.
Let $J\in\mathbb{I}$ be such that $|J_{n}|=(n+1)^{2}$ for all $n<\omega$, and fix $I\in\mathbb{I}$. Set $I_{n}^{\prime}:=\bigcup_{k\in J_{n}}I_{k}$, so $I^{\prime}\in\mathbb{I}$. Since $D$ is $\mathbb{I}$-dominating, we can find an $I^{*}\in D$ such that, for all but finitely many $n<\omega$, $I^{*}$ contains at least two intervals from $I^{\prime}$. Therefore, there is some $\varphi^{*}\in\mathcal{S}({}^{I^{*}}2,\operatorname{id})$ such that $X\subseteq H_{\varphi^{*}}$. It is enough to construct a $\varphi\in\mathcal{S}({}^{I}2,\operatorname{id})$ such that $H_{\varphi^{*}}\subseteq H_{\varphi}$.
We first show that, for all but finitely many $n<\omega$ and for any $k<\omega$, if $I_{k}\subseteq I^{*}_{n}$ then $k\geq n(n+1)$: For any $\ell<\omega$, $I^{\prime}_{\ell}$ is composed of $(\ell+1)^{2}$-many intervals from $I$, so $I_{k}\subseteq I^{\prime}_{\ell}$ iff $\frac{\ell(\ell+1)(2\ell+1)}{6}\leq k<\frac{(\ell+1)(\ell+2)(2\ell+3)}{6}$. Now, for all but finitely many $n$, we have that whenever $I^{\prime}_{\ell}$ intersects $I^{*}_{n}$, $\ell\geq n$ and $\ell(\ell+1)<\frac{(\ell-1)\ell(2\ell-1)}{6}$. Therefore, for any $I_{k}\subseteq I^{*}_{n}$, $n(n+1)\leq\ell^{\prime}(\ell^{\prime}+1)<\frac{(\ell^{\prime}-1)\ell^{\prime}(2\ell^{\prime}-1)}{6}\leq k$, where $\ell^{\prime}$ is the smallest $\ell$ such that $I^{\prime}_{\ell}\cap I^{*}_{n}\neq\emptyset$.
Fix an $n$ as above. Whenever $I_{k}\subseteq I^{*}_{n}$, define $\varphi(k):=\{t{\restriction}I_{k}\colon t\in\varphi^{*}(n)\}$, so $|\varphi(k)|\leq|\varphi^{*}(n)|\leq n\leq n(n+1)\leq k$; and whenever $I_{k}$ intersects both $I^{*}_{n}$ and $I^{*}_{n+1}$, define
$\varphi(k):=\{s{\restriction}(I_{k}\cap I_{n}^{*})\cup t{\restriction}(I_{k}\cap I_{n+1}^{*})\colon s\in\varphi^{*}(n),t\in\varphi^{*}(n+1)\},$
so $|\varphi(k)|\leq|\varphi^{*}(n)|\cdot|\varphi^{*}(n+1)|\leq n(n+1)\leq k$.
So far, we have defined $\varphi(k)$ for all but finitely many $k<\omega$, so for the remaining ones set $\varphi(k):=\emptyset$. Therefore, $\varphi\in\mathcal{S}({}^{I}2,\operatorname{id})$.
Lastly, it is clear that $H_{\varphi^{*}}\subseteq H_{\varphi}$, as required. ∎
Thanks to the above, we have the following convenient Tukey connection.
Theorem 4.4.
For any dominating family $D\subseteq\omega^{\omega}$, we have $\mathrm{Cv}_{\mathcal{N}\!\mathcal{A}}\leq_{\mathrm{T}}\prod_{b\in D}\mathrm{Lc}(b,\operatorname{id})$.
Proof.
Without loss of generality, we may assume that there is some $\mathbb{I}$-dominating family $D_{0}$ such that, for each $b\in D$, there is some $I\in D_{0}$ such that $b={}^{I}2$, i.e. $b(n)={}^{I_{n}}2$ for all $n<\omega$.
We define $\Psi_{-}\colon{}^{\omega}2\to\prod_{I\in D_{0}}{}^{I}2$ and $\Psi_{+}\colon\prod_{I\in D_{0}}\mathcal{S}({}^{I}2,\operatorname{id})\to\mathcal{N}\!\mathcal{A}$ as follows: For $x\in{}^{\omega}2$ and $I\in D_{0}$, set $x_{I}:=\langle x{\restriction}I_{n}\colon n\in\omega\rangle$ and define $\Psi_{-}(x):=\langle x_{I}\colon I\in D_{0}\rangle$; for $\bar{\varphi}=\langle\varphi_{I}\colon I\in D_{0}\rangle\in\prod_{I\in D_{0}}\mathcal{S}({}^{I}2,\operatorname{id})$ define
$\Psi_{+}(\bar{\varphi}):=\bigcap_{I\in D_{0}}\{x\in{}^{\omega}2\colon\forall^{\infty}n\in\omega\colon x{\restriction}I_{n}\in\varphi_{I}(n)\}.$
Notice that $\Psi_{+}(\bar{\varphi})\in\mathcal{N}\!\mathcal{A}$ by Theorem 4.3. It is clear that, for any $x\in{}^{\omega}2$ and $\bar{\varphi}=\langle\varphi_{I}\colon I\in D_{0}\rangle\in\prod_{I\in D_{0}}\mathcal{S}({}^{I}2,\operatorname{id})$, $\Psi_{-}(x)\sqsubset^{\otimes}\bar{\varphi}$ implies $x\in\Psi_{+}(\bar{\varphi})$. ∎
We are ready to prove our consistency results. First, we review some terminology.
Definition 4.5.
Define the following forcing notions
  1. (1)
    Fix $b$ and $h$ as in 4.1. Define $\mathbb{LOC}_{b,h}$ as the poset whose conditions are pairs $p=(\varphi_{p},n_{p})$ where $\varphi_{p}\in\mathcal{S}(b,h)$ and $n_{p}<\omega$ such that, for some $m_{p}<\omega$, $|\varphi_{p}(i)|\leq m_{p}$ for all $i<\omega$. The order is defined by $q\leq p$ iff $n_{p}\leq n_{q}$, $\varphi_{q}(i)=\varphi_{p}(i)$ for all $i<n_{p}$ and $\forall i<\omega\colon\varphi_{p}(i)\subseteq\varphi_{q}(i)$.
  2. (2)
    Hechler forcing is defined by $\mathbb{D}=\omega^{<\omega}\times{}^{\omega}\omega$, ordered by $(t,g)\leq(s,f)$ if $s\subseteq t$, $f\leq g$ and $f(i)\leq t(i)$ for all $i\in|t|\smallsetminus|s|$. This forcing is used to increase $\mathfrak{b}$. Recall that $\mathbb{D}$ is $\sigma$-centered.
  3. (3)
    For an infinite cardinal $\theta$, $\mathrm{Fn}_{<\theta}(A,B)$ denotes the poset of partial functions from $A$ into $B$ of size $\theta$, ordered by $\supseteq$.
Theorem 4.6.
Let $\theta\leq\lambda$ and $\nu$ be uncountable cardinals such that $\theta$ is regular, $\lambda^{\aleph_{0}}=\lambda$ and $\mathfrak{d}_{\theta}=\nu$. [^4] Notice that no inequality is assumed between $\nu$ and $\lambda$. Then there is a ccc poset forcing
$\operatorname{add}(\mathcal{N})=\operatorname{cof}(\mathcal{N})=\theta\leq\operatorname{cov}(\mathcal{N}\!\mathcal{A})\leq\nu\text{ and }\mathfrak{c}=\lambda.$
In particular, it is consistent with $\mathrm{ZFC}$ that $\operatorname{cov}(\mathcal{N}\!\mathcal{A})<\lambda$ (starting with $\nu<\lambda$).
Here, $\mathfrak{d}_{\theta}$ is the canonical dominating number of $^{\theta}\theta$, which coincides with $\mathfrak{d}(\langle\theta,\theta,\leq\rangle^{\theta})$. It is easy to force $\mathfrak{d}_{\theta}=\nu$, e.g. with $\mathrm{Fn}_{<\theta}(\nu,2)$ (by assuming $\theta^{<\theta}=\theta$ and $\nu^{\theta}=\nu$ in the ground model).
Proof.
Perform a FS iteration $\mathbb{P}=\langle\mathbb{P}_{\xi},\dot{\mathbb{Q}}_{\xi}\colon\xi<\lambda\theta\rangle$ ($\lambda\theta$ as product of ordinal numbers) where $\mathbb{Q}_{\xi}$ is a $\mathbb{P}_{\xi}$-name of $\mathbb{D}\ast\mathbb{LOC}_{\dot{d}_{\xi},\operatorname{id}}$ where $\dot{d}_{\xi}$ is the name of the dominating real over $V_{\xi}:=V^{\mathbb{P}_{\xi}}$ added by $\mathbb{D}$. Using the cofinaly-many Cohen and dominating reals $\langle\dot{d}_{\lambda\rho}\colon\rho<\theta\rangle$ added along the iteration and $\operatorname{cf}(\lambda\theta)=\theta$, we obtain $\mathfrak{b}=\operatorname{non}(\mathcal{M})=\operatorname{cov}(\mathcal{M})=\mathfrak{d}=\theta$. which implies $\operatorname{add}(\mathcal{M})=\operatorname{cof}(\mathcal{M})=\theta$. Even more, we obtain $\omega^{\omega}\cong_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}}\cong_{\mathrm{T}}\theta$.
In the final generic extension $V_{\lambda\theta}$, it is clear that $D:=\{d_{\lambda\rho}\colon\rho<\theta\}$ is $\leq^{*}$-increasing and dominating in $^{\omega}\omega$. Denote $d^{\prime}_{\rho}:=d_{\lambda\rho}$ for $\rho<\theta$. We show that $\mathrm{Lc}(d^{\prime}_{\rho},\operatorname{id})\cong_{\mathrm{T}}\theta$. On the one hand, $\theta\cong_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}}\leq_{\mathrm{T}}\mathrm{Lc}(d^{\prime}_{\rho},\operatorname{id})$. For the converse, define $F\colon\prod d^{\prime}_{\rho}\to\theta$ such that, for $x\in\prod d^{\prime}_{\rho}$, $F(x)$ is some ordinal $\eta>\rho$ such that $x\in V_{\lambda\eta}$; and define $F^{\prime}\colon\theta\to\mathcal{S}(d^{\prime}_{\rho},\operatorname{id})$ such that $F^{\prime}(\varrho)$ is the $\mathbb{LOC}_{d^{\prime}_{\varrho},\operatorname{id}}$-generic real added by $\mathbb{Q}_{\lambda\varrho}$ when $\varrho\geq\rho$, otherwise $F^{\prime}(\varrho):=F^{\prime}(\rho)$. It is clear that $(F,F^{\prime})$ is the desired Tukey connection.
Since $\mathrm{Lc}(d^{\prime}_{\rho},\operatorname{id})\cong_{\mathrm{T}}\theta$, we get $\mathrm{minLc}=\mathrm{supLc}=\theta$, so by Theorem 4.2 $\operatorname{add}(\mathcal{N})=\operatorname{cof}(\mathcal{N})=\theta$. On the other hand, given that $\mathbb{P}$ is ccc, the equality $\mathfrak{d}_{\theta}=\nu$ is preserved (see e.g. [CM25, Lem. 6.6]). Now, by Theorem 4.4,
$\mathrm{Cv}_{\mathcal{N}\!\mathcal{A}}\leq_{\mathrm{T}}\prod_{b\in D}\mathrm{Lc}(b,\operatorname{id})\cong_{\mathrm{T}}\langle\theta,\leq\rangle^{\theta},$
so $\operatorname{cov}(\mathcal{N}\!\mathcal{A})\leq\mathfrak{d}_{\theta}=\nu$. ∎
Finally, we show that $\operatorname{cof}(\mathcal{M}\mathcal{A})<\operatorname{non}(\mathcal{SN})$ is valid in our model from [CMR25, Thm. 4.2].
Theorem 4.7.
Let $\theta\leq\lambda$ and $\nu$ be uncountable cardinals such that $\theta$ is regular, $\mathfrak{d}_{\theta}=\nu$ and $\lambda^{\aleph_{0}}=\lambda$. Then there is a ccc poset forcing
$\operatorname{cov}(\mathcal{N})=\aleph_{1}\leq\operatorname{add}(\mathcal{M})=\operatorname{cof}(\mathcal{M})=\theta\leq\operatorname{non}(\mathcal{N})=\mathfrak{c}=\lambda\text{ and }\operatorname{cof}(\mathcal{M}\mathcal{A})\leq\nu.$
In particular, it is consistent with $\mathrm{ZFC}$ that $\operatorname{cof}(\mathcal{M}\mathcal{A})<\operatorname{non}(\mathcal{N})$.
Proof.
Let $\mathbb{P}$ be the FS iteration of ccc posets constructed in the proof of [CMR25, Thm. 4.2]. It remains to show that $\mathbb{P}$ forces $\operatorname{cof}(\mathcal{M}\mathcal{A})\leq\nu$. Since $\mathbb{P}$ has the ccc, the identity $\mathfrak{d}_{\theta}=\nu$ is preserved in the generic extension. Now, in the generic extension, by 3.18, $\operatorname{cof}(\mathcal{M}\mathcal{A})\leq\operatorname{cov}\left(([\theta]^{<\theta})^{\theta}\right)=\mathfrak{d}_{\theta}=\nu$. ∎
We still do not know how to control the values of $\operatorname{cov}(\mathcal{N}\!\mathcal{A})$, $\operatorname{cov}(\mathcal{M}\mathcal{A})$ and $\operatorname{cof}(\mathcal{M}\mathcal{A})$ in the previous models. In the case of the cofinality, we may need to develop some framework for lower bounds as we did in [CM25] for the Yorioka ideals. In particular, we ask:
Question 4.8.
Is it consistent with $\mathrm{ZFC}$ that $\operatorname{cof}(\mathcal{M}\mathcal{A})>\mathfrak{c}$?
Concerning $\mathcal{N}\!\mathcal{A}$ we may ask:
Question 4.9.
Does $\mathrm{ZFC}$ proves an inequality between $\operatorname{cof}(\mathcal{N}\!\mathcal{A})$ and $\mathfrak{c}$?
We have some ideas to construct a directed scheme for $\mathcal{N}\!\mathcal{A}$ to prove the consistency of $\operatorname{cof}(\mathcal{N}\!\mathcal{A})<\mathfrak{c}$, but we will develop them in the second part of this work.